### v2

# Symmetry Breaking Dynamics in Quantum Many-Body Systems

Abstract
Entanglement asymmetry has emerged as a powerful tool for characterizing symmetry breaking in quantum many-body systems. In this Letter, we explore how symmetry is dynamically broken through the lens of entanglement asymmetry in two distinct scenarios: a non-symmetric random quantum circuit and a non-symmetric Hamiltonian quench, with a particular focus on U(1) symmetry. In the former case, the symmetry is initially broken and subsequently restored, whereas in the latter case, symmetry remains broken in the subsystem at late times, consistent with the principles of quantum thermalization. Notably, the growth of entanglement asymmetry exhibits unexpected overshooting behavior at early times in both contexts, contrasting with the behavior of charge variance. We also consider dynamics of non-symmetric initial states under the symmetry-breaking evolution. Due to the competition of symmetry-breaking in both the initial state and Hamiltonian, the early-time entanglement asymmetry can increase and decrease, while quantum Mpemba effects remain evident despite the weak symmetry-breaking in both settings.
Introduction.— Symmetry breaking is a ubiquitous phenomenon across all branches of physics. A well-known example is the Higgs mechanism [1] in particle physics, where the vacuum state of the universe causes different particles to acquire mass, spontaneously breaking the electroweak symmetry. This type of symmetry breaking, which occurs without external influences, is referred to as spontaneous symmetry breaking. In contrast, a symmetry can also be explicitly broken when the Hamiltonian describing the system directly breaks the symmetry. How symmetry breaks dynamically in this case is an interesting fundamental question to explore.
Symmetry properties are also closely related to the concept of quantum thermalization [2, 3, 4, 5, 6] for generic quantum many-body systems. In general, when a closed quantum system evolves with a chaotic Hamiltonian, the reduced density matrix of a small subsystem $a$ thermalizes to the equilibrium finite-temperature state: $\rho_{a}\propto e^{-\beta\hat{H}_{a}}$ where $\hat{H}_{a}$ is the Hamiltonian of the subsystem. Symmetry is restored at later times for symmetric Hamiltonian $\hat{H}_{a}$, since $[\hat{Q}_{a},\rho_{a}]=0$ where $\hat{Q}_{a}$ represents the corresponding symmetry generator. However, if $\hat{H}_{a}$ does not respect the symmetry, the reduced density matrix $\rho_{a}$ at late times is non-commuting with $\hat{Q}_{a}$. In this case, symmetry breaking persists even if the system begins in a symmetric state.
Despite the richness of the late-time behavior, early-time dynamics have also garnered significant attention in recent years, particularly due to the novel Mpemba effect [7], which demonstrates that hot water freezes faster than cold water. Both classical and quantum versions of the Mpemba effect have been widely explored in various systems [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Recently, quantum Mpemba effect (QME) is reported in quantum integrable systems and chaotic systems [24, 25, 26]. Notably, U(1)-symmetry restoration occurs more rapidly for more asymmetric initial states under the U(1)-symmetric Hamiltonian quench [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. This finding was subsequently explored in various other settings [38, 39, 40, 41, 42, 43, 44, 45, 46, 47] and experimentally realized on a trapped-ion quantum simulator [48].
Previous studies [24, 25] have primarily focused on characterizing symmetry restoration when an asymmetric initial state evolves under a symmetric Hamiltonian or random circuit. In contrast, this Letter examines the dynamical aspects of symmetry breaking, exploring the behavior of symmetric and asymmetric initial states under non-symmetric evolution [49]. In addition, due to experimental limitations, symmetric evolutions are often affected by noises and defects, resulting in non-symmetric contributions as well. In such cases, can symmetry restoration still occur, or does symmetry breaking become more pronounced over time? Additionally, how does QME behave in the presence of symmetry-breaking interactions? Addressing these questions offers a more comprehensive understanding of symmetry and symmetry breaking in quantum many-body systems.
Figure 1: (a) Schematic illustration of a non-symmetric random circuit with 6 qubits. Gates are arranged in the even-odd brick-wall pattern. The blue and red rectangles represent U(1)-symmetric and random Haar gates, respectively. The basis for the U(1)-symmetric gate is listed in the following order: $|00\rangle$, $|01\rangle$, $|10\rangle$ and $|11\rangle$. (b) The circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]$, as a function of time with the antiferromagnetic initial state at different values of $P_{\text{Haar}}$. (c) The peak value, $\mathbb{E}[\Delta S_{L/4}]_{max}$, as a function of $P_{\text{Haar}}$. All three curves follow a power law $y=ax^{b}$. F: Ferromagnetic state ($a=1.4$, $b=0.4$); DW: Domain Wall state ($a=2.7$, $b=0.8$); AF: Antiferromagnetic state ($a=1.9$, $b=0.9$).
In this Letter, we investigate and compare the dynamics of symmetry breaking with two distinct models: a non-symmetric random circuit [50] and a non-symmetric Hamiltonian evolution, each with different symmetric and asymmetric initial states. To characterize the extent of symmetry breaking in subsystem $a$, we employ the metric of entanglement asymmetry (EA) [24], which has been extensively utilized as a measure of symmetry breaking in quantum field theories [51, 52, 53] and out-of-equilibrium many-body systems [39, 41, 54]. EA is defined as
$\Delta S_{a}=S(\rho_{a,Q})-S(\rho_{a}).$ (1)
Here, $S(\rho_{a})$ denotes the standard Von Neumann entropy of subsystem $a$, and $\rho_{a,Q}=\sum_{q\in\mathbb{Z}}\Pi_{q}\rho_{a}\Pi_{q}$ where $\hat{Q}_{a}=\sum_{i\in a}\sigma_{i}^{z}$ in case of U(1) symmetry and $\Pi_{q}$ is the projector onto eigenspace of $\hat{Q}_{a}$ with charge $q$. Consequently, $\rho_{a,Q}$ is block diagonal in the eigenbasis of $\hat{Q}_{a}$. The EA satisfies two key properties: (1) $\Delta S_{a}\geq 0$ since the EA is defined as the relative entropy between $\rho_{a,Q}$ and $\rho_{a}$. (2) $\Delta S_{a}=0$ if and only if $\rho_{a,Q}=\rho_{a}$. In random circuit settings, $\mathbb{E}[\Delta S_{a}]$ is employed as the circuit-averaged value of $\Delta S_{a}$. Note that the symmetry for subsystem mixed states investigated here corresponds to the weak symmetry in Refs. [55, 56]. In parallel with the analysis of EA, we also compute the charge variance (CV) $\sigma^{2}_{Q}=\langle\hat{Q}^{2}\rangle-\langle\hat{Q}\rangle^{2}$, where $\hat{Q}=\sum_{i=1}^{L}\sigma_{i}^{z}$. This quantity serves as a measure of charge fluctuations within the system, offering a complementary perspective on the dynamics of symmetry breaking.
In the case of non-symmetric random circuits, we show that U(1) symmetry for a small subsystem can still be restored regardless of whether the initial state is U(1)-symmetric or asymmetric. As a result, EA exhibits overshooting at early times, i.e. there is a peak of EA at an early time that is much larger than the saturating EA value at late times. Additionally, QME appears at early times unless all U(1)-symmetric gates are replaced by random Haar gates, where EA dynamics are exactly the same for different U(1)-asymmetric initial states.
For a non-symmetric Hamiltonian evolution, we find that U(1) symmetry can not be restored in a subsystem. This behavior can be explained by the late-time reduced density matrix relaxing to the form $e^{-\beta\hat{H}_{a}}$, where $\hat{H}_{a}$ explicitly includes symmetry-breaking terms. In this scenario, the EA still shows nontrivial overshooting at early times, contrasting to other symmetry-breaking measures such as charge variance. Furthermore, early-time EA dynamics from asymmetric initial states display distinct behaviors due to the competition between two symmetry-breaking contributions related to the initial state and the system Hamiltonian. Moreover, the QME originated from symmetric evolution disappears when the strength of symmetry breaking in the evolution exceeds some thresholds.
Setup.— To characterize symmetry breaking in these systems, we consider three initial states: the ferromagnetic state $|000...0\rangle$, the antiferromagnetic state $|0101..1\rangle$ and the domain-wall state $|000..111\rangle$, where the domain wall is positioned at the center of the chain. To incorporate the effect of symmetry breaking in the initial state, we introduce tilted ferromagnetic states, as discussed in earlier studies [24, 25]. The state is defined as
$|\psi_{i}(\theta)\rangle=e^{-i\frac{\theta}{2}\sum_{j}\sigma_{j}^{y}}|000...0\rangle$ (2)
where $\sigma_{j}^{y}$ is the Pauli matrix in $y$-direction acting on the $j$-th qubit, and $\theta$ is a tuning parameter that controls the strength of symmetry breaking in the initial state. When $\theta=0$, Eq. (2) is U(1)-symmetric, resulting in a vanishing EA. As $\theta$ increases, the EA grows, reaching its maximum value at $\theta=\pi/2$. The tilted antiferromagnetic and tilted domain wall states are constructed in a similar manner.
The random circuit in Fig. 1 (a) consists of two-qubit random U(1)-symmetric gates and random Haar gates, arranged in a brick-wall fashion. The exact matrix form of U(1)-symmetric gate is illustrated in Fig. 1 (a), where each block is randomly drawn from the Haar measure [57, 58, 59]. The effect of symmetry breaking depends on the density (doped probability) of random Haar gates without U(1) symmetry, denoted as $P_{\text{Haar}}$. The time unit in the circuit is defined by the application of two consecutive layers of gates. The initial state $|\psi_{i}(\theta)\rangle$ evolves under the random unitary dynamics and $\mathbb{E}[\Delta S_{a}]$ is computed by averaging $\Delta S_{a}$ over $5000$ circuit configurations.
We also investigate Hamiltonian dynamics where the state $|\psi_{i}(\theta)\rangle$ undergoes unitary evolution given by $e^{-iHt}|\psi_{i}(\theta)\rangle$, and the Hamiltonian is defined as
$H=$ $-\frac{1}{4}\sum_{j=1}^{L}\Big{[}\sigma_{j}^{x}\sigma_{j+1}^{x}+\gamma\sigma_{j}^{y}\sigma_{j+1}^{y}+\Delta_{1}\sigma_{j}^{z}\sigma_{j+1}^{z}\Big{]}$
$-\Delta_{2}\sum_{j=1}^{L}\Big{[}\sigma_{j}^{x}\sigma_{j+2}^{x}+\sigma_{j}^{y}\sigma_{j+2}^{y}+\sigma_{j}^{z}\sigma_{j+2}^{z}\Big{]}.$
Here, $\Delta_{1}$ and $\Delta_{2}$ are the coefficients for nearest-neighbor and next-nearest-neighbor interactions, respectively. $\Delta_{2}$ introduces non-integrability, and $\gamma$ controls the strength of symmetry breaking. Periodic boundary conditions are imposed in both contexts.
Figure 2: The circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]$, as a function of time for different values of $P_{\text{Haar}}$. Blue: U(1) EA. Green: $Z_{2}$ EA. Panels (a)-(d) correspond to different values of $P_{\text{Haar}}$ (a) $P_{\text{Haar}}=0$, (b) $P_{\text{Haar}}=0.3$, (c) $P_{\text{Haar}}=0.7$, and (d) $P_{\text{Haar}}=1$, respectively.
U(1)-Symmetric (Asymmetric) States with U(1) Non-Symmetric Random Circuit.— All numerical simulations are performed using the TensorCircuit-NG package [60]. The circuit under investigation consists of 16 qubits. We evaluate the EA at different doping probabilities of random Haar gates, using an antiferromagnetic initial state. We observe that, at later times, all EAs approach zero, as illustrated in Fig. 1 (b). This behavior can be understood in the context of quantum thermalization and information scrambling [61, 62, 63, 64], where the reduced density matrix of the subsystem is a fully mixed state for the random circuit cases, as long as the subsystem size does not exceed half of the total system size. Additionally, all EAs display a peak at some early time steps. For all probabilities chosen in Fig. 1 (b), all EAs reach their maximum after only a few layers of unitaries. The rate of symmetry restoration also depends on the initial state. In the supplementary material (SM), we find that symmetry restoration occurs more quickly for antiferromagnetic or domain wall states than for ferromagnetic states, due to the larger Hilbert space sector of the initial states in the former cases. In Fig. 1 (c), we reveal that the peak of the circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]_{max}$ follows a power-law with respect to $P_{\text{Haar}}$.
Next, we examine the dynamics from U(1)-asymmetric initial states, i.e. a tilted ferromagnetic state. We compute the EA for both U(1) symmetry with $\hat{Q_{a}}=\sum_{i\in a}\sigma_{i}^{z}$ and $Z_{2}$ symmetry with $\hat{Q_{a}}=\prod_{i\in a}\sigma_{i}^{z}$. As depicted in Fig. 2 (a), for $P_{\text{Haar}}=0$, we clearly notice the emergence of QME in U(1) case. Surprisingly, we also find that the QME appears in the $Z_{2}$ probe, which does not contradict previous study [25] suggesting the absence of QME in $Z_{2}$-symmetric circuits. Even though U(1)-symmetric gates are also $Z_{2}$ symmetric, there is no off-diagonal coupling between $|00\rangle$ and $|11\rangle$, leading to different thermalization rates between two $Z_{2}$ charge sectors ($Q_{a}=\pm 1$), and thus resulting in QME.
Figure 3: EA as a function of time with (a) ferromagnetic and (b) antiferromagnetic states for different values of $\gamma$ under $H_{1}$. The insets show the peak of EA at different values of $\gamma$. From bottom to top: $\gamma=0.8,0.7,0.6,0.5,0.4,0.3,0.2,0.1$. Panels (c) and (d) show the peak value of EA, $(\Delta S_{L/3})_{max}$, and the ratio of the late-time EA, $\Delta S_{L/3}^{\infty}$, to $(\Delta S_{L/3})_{max}$ as a function of $1-\gamma$ for various initial states under $H_{1}$. GS denotes the value of EA calculated from the ground state of $H_{1}$.
As we replace a portion of U(1)-symmetric gates with random Haar gates, QME remains evident with a finite number of random Haar gates. However, when the circuit consists entirely of random Haar gates, all charge sectors thermalize at the same rate after circuit averaging, and QME disappears.
U(1)-Symmetric Initial States with U(1) Non-Symmetric Hamiltonian.— Here, we investigate the dynamics of symmetry breaking under an integrable Hamiltonian $H_{1}$ with $\Delta_{1}=0.4$ and $\Delta_{2}=0$, and a non-integrable Hamiltonian $H_{2}$ with $\Delta_{1}=0.4$ and $\Delta_{2}=0.05$, with system size 12 sites. As revealed in Fig. 3 (a) and (b), we calculate EA for various Hamiltonian symmetry-breaking values $\gamma$ and observe that EAs also exhibit peaks at early times that are much larger than steady values. Furthermore, the peak value of the EA, $(\Delta S_{L/3})_{max}$, is found to be correlated with the strength of symmetry breaking, $1-\gamma$, for different symmetric initial states as shown in Fig. 3 (c) where EA of the ground state of $H_{1}$ follows the same trend. Notably, the peak heights nearly coincide between the ferromagnetic and domain wall states, as the early-time peak primarily depends on the local configurations of the initial state.
Figure 4: EA dynamics for (a) tilted ferromagnetic states and (b) tilted antiferromagnetic states with varying $\gamma$. The blue curves correspond to $\theta=0.2\pi$, and the red curves represent $\theta=0.5\pi$. Panels (c) and (d) depict schematic 2-dimensional phase diagrams illustrating the dependence of early-time EA dynamics on $\theta$ and $1-\gamma$ for ferromagnetic and antiferromagnetic states, respectively. When the parameter is in the red region, EA can exceed the initial value, while in the blue region, EA firstly decreases and never grows higher than the initial value. All black dots are obtained through numerical simulation. All calculations are based on $H_{1}$.
By analyzing Fig. 3, we identify that the late-time EA, denoted as $\Delta S_{L/3}^{\infty}$, oscillates and does not approach zero. This is because the reduced density matrix of subsystem $a$ evolves towards a canonical ensemble $e^{-\beta\hat{H_{a}}}$, where $\hat{H_{a}}$ has the same form as $\hat{H}$ in Eq. (Symmetry Breaking Dynamics in Quantum Many-Body Systems), but acts solely on subsystem $a$. Since $\hat{H_{a}}$ includes symmetry breaking terms, $[\rho_{a},\hat{Q_{a}}]\neq 0$, leading to a non-vanishing EA at long times. In Fig. 3 (d), we calculate the ratio of $\Delta S_{L/3}^{\infty}$ to $(\Delta S_{L/3})_{max}$ with varying $\gamma$ for different initial states. The late-time EA, $\Delta S_{L/3}^{\infty}$, is obtained by averaging $\Delta S_{L/3}$ over $2000$ time points between $t_{1}=2000$ and $t_{2}=40000$. The results further confirm the overshooting behavior as the late-time saturating EA value is much lower than the peak value at the early time. On the contrary, the CV dynamics in this setting shows no evident overshooting pattern but instead directly grows to the saturating values (see SM).
U(1)-Asymmetric Initial States with U(1) Non-Symmetric Hamiltonian.— Next we investigate the behavior of EA with U(1)-asymmetric initial states under $H_{1}$. Now, the time evolution of EA depends on both symmetry-breaking parameters, $\theta$ and $\gamma$. $\theta$ describes the symmetry breaking in the initial state and determines the initial value of EA. $\gamma$, on the other hand, serves as a measure of symmetry breaking in the Hamiltonian, which influences the evolution of the state. The interplay between these two parameters results in distinct behaviors in the time dependence of EA. This is illustrated in the schematic 2-dimensional phase diagrams with varying $\theta$ and $\gamma$ in Fig. 4 (c) and (d). The colors highlight the tendency in EA at early times. Blue regime indicates that $\Delta S_{L/4}(t)$ never exceeds its initial value for early times, while red regime corresponds to the situations where EA can grow larger than its initial value at early times. It is clearly reflected in Figs. 4 (a), the initial growth of EA at $\theta=0.2\pi$ and $\gamma=0.8$, $0.6$ aligns with the red region shown in the phase diagram. Two key observations are made from the phase diagram: (1) For a fixed $\gamma$, EA grows with weaker asymmetric effects (small $\theta$) in the initial states or for a fixed $\theta$, EA increases with stronger symmetry breaking effects (large $1-\gamma$) in the Hamiltonian. (2) The threshold for $\gamma$ at which EA begins to rise varies with different initial states. As illustrated in Fig. 4 (d), the range in which EA exceeds the initial value is very limited with an antiferromagnetic initial state. In other words, the early-time behavior of EA serves as a witness to compare the symmetry-breaking strength hosting in the quantum state and the Hamiltonian. When the Hamiltonian is more asymmetric than the initial state, EA will first increase, often resulting in the overshooting behavior in the early time, as indicated by the red regime in the phase diagram.
Another key feature of the early-time dynamics is the emergence of QME, as shown in Fig. 4 (a) for the symmetric case $\gamma=1$. The origin of this QME lies in the relative small $ZZ$ term and gapless nature in the Hamiltonian [42]. Notably, QME persists for ferromagnetic (antiferromagnetic) states when $0.8\leq\gamma\leq 1$ ($0.4\leq\gamma\leq 1$). The robustness of QME against weak symmetry-breaking is a general feature in quantum many-body systems when the Hamiltonian is non-symmetric. We also report relevant results for non-integrable Hamiltonian $H_{2}$ in the SM, and the results remain qualitatively consistent with cases of $H_{1}$, demonstrating the universal applicability of conclusions in this Letter for Hamiltonian evolutions.
Our simulation on another symmetry-breaking measure, charge variance, reveals that QME can also emerge for CV with initial tilted ferromagnetic states, but only in cases of non-symmetric evolution. The reversed monotonicity of CV with respect to $\theta$ can persist even at late times. Table. 1 summarizes the early- and late-time behavior of EA and CV for different initial states. The distinction shows the richness in characterizing symmetry breaking strength and patterns.
Ferromagnetic Domain Wall Antiferromagnetic
$EA\ (early\ time)$ crossing for small $1-\gamma$ crossing for small $1-\gamma$ crossing for small $1-\gamma$
$CV\ (early\ time)$ crossing for $\gamma\neq 1$ no crossing no crossing
$EA\ (late\ time)$
$CV\ (late\ time)$
Table 1: The early- and late-time behavior of EA and CV under the evolution of $H_{1}$ or $H_{2}$ ($0.5\leq\gamma\leq 1$). The crossing observed at early times signals the occurrence of the QME. The right-up (right-down) arrow indicates that the late-time value is increasing (decreasing) with increasing tilted angle $\theta$.
Conclusions and discussions.— In this Letter, we present a comprehensive study of subsystem symmetry breaking within two frameworks: a non-symmetric random circuit and a non-symmetric Hamiltonian evolution. Our simulation reveals that U(1) symmetry is always restored in the non-symmetric random circuit case, regardless of the initial states or the density of symmetry-breaking random Haar gates $P_{\text{Haar}}$. On the contrary, subsystem U(1) symmetry remains broken in the case of a U(1) non-symmetric Hamiltonian.
In addition to the late-time results, the early-time dynamics of EA shows a universal and surprising feature of overshooting. Specifically, the initial growth of EA can reach a peak significantly higher than its late-time steady value. This behavior is unexpected and is distinct from the growth of entanglement or charge variance, another measure of symmetry-breaking, where the value increases monotonically to its saturating level without any evident overshooting. Furthermore, for asymmetric initial states evolved under non-symmetric Hamiltonians, the distinct and rich early-time dynamics of EA (increase versus decrease) allow for a direct comparison of the symmetry-breaking extent in both the state and the Hamiltonian.
There are several promising directions for further exploration. For instance, studying the dynamics of symmetry breaking in a non-unitary random circuit with mid-circuit measurements [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81], could offer valuable insights. Additionally, examining the effect of symmetry breaking in Hamiltonians that avoid thermalization such as many-body localization systems [45, 82, 83, 84, 85, 86, 87, 88, 89, 90] can provide a more unified picture on the understanding of symmetry-breaking dynamics.
Acknowledgement.— HY is supported by the International Young Scientist Fellowship of Institute of Physics Chinese Academy of Sciences (No.202407). ZXL is supported by the National Natural Science Foundation of China under the Grant No.12347107 and Grant No.12474146. SXZ is supported by a start-up grant at IOP-CAS.
Data availability. Numerical data for this manuscript are publicly accessible in Ref. [91].
Supplemental Material for “Symmetry Breaking Dynamics in Quantum Many-Body Systems”
February 8, 2025
I More numerical results for U(1) non-symmetric quantum circuits
I.1 Dynamics of EA with other U(1)-symmetric initial states
In this section, we present numerical results of the dynamics of entanglement asymmetry with different initial states, using the same setup as described in the main text. Similar to the behavior observed in Fig. 1 (b), EA for the other initial states in Fig. S1 also exhibits an initial peak, which then gradually decays to zero, indicating the restoration of U(1) symmetry. It is most apparent from the case of ferromagnetic states that a smaller density of random Haar gates results in a longer time of symmetry restoration in the subsystem. In addition, the thermalization process occurs more rapidly in domain wall and antiferromagnetic states than ferromagnetic states. This is because, in the limit of symmetry-preserving circuits, the Hilbert space accessible to the antiferromagnetic states is much larger than the ferromagnetic cases.
Figure S1: The dynamics of circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]$ for different initial states with $L=16$. The varying color intensity represents different densities of random Haar gates. Left: Ferromagnetic state. Right: Domain Wall state.
I.2 Dynamics of EA with U(1)-asymmetric initial states at different $P_{\text{Haar}}$
In addition to the $P_{\text{Haar}}$ values used in the main text, we show the behavior of entanglement asymmetry with tilted ferromagnetic state for other $P_{\text{Haar}}$ values in Fig. S2, and find that QME is indeed present except when the symmetry breaking reaches the maximum $P_{\text{Haar}}=1$. This phenomenon also holds true for tilted domain wall state, as demonstrated in Fig. S3. On the contrary, as shown in Fig. S4, QME is absent for tilted antiferromagnetic states across all values of $P_{\text{Haar}}$. Therefore, the emergence of QME is specific to certain initial states and is robust against the effects of random Haar gates on the circuit evolution.
Figure S2: The dynamics of circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]$, for tilted ferromagnetic initial states is examined for various values of $P_{\text{Haar}}$ with $L=16$. Blue: U(1) EA. Green: $Z_{2}$ EA. Panels (a)-(d) correspond to the following values of $P_{\text{Haar}}$ (a) $P_{\text{Haar}}=0.1$, (b) $P_{\text{Haar}}=0.2$, (c) $P_{\text{Haar}}=0.5$, and (d) $P_{\text{Haar}}=0.9$.
Figure S3: The dynamics of circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]$, for tilted domain wall initial states is examined for various values of $P_{\text{Haar}}$ with $L=16$. Blue: U(1) EA. Green: $Z_{2}$ EA. Panels (a)-(d) correspond to the following values of $P_{\text{Haar}}$ (a) $P_{\text{Haar}}=0$, (b) $P_{\text{Haar}}=0.3$, (c) $P_{\text{Haar}}=0.7$, and (d) $P_{\text{Haar}}=1$.
Figure S4: The dynamics of circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]$, for tilted antiferromagnetic initial states is examined for various values of $P_{\text{Haar}}$ with $L=16$. Blue: U(1) EA. Green: $Z_{2}$ EA. Panels (a)-(d) correspond to the following values of $P_{\text{Haar}}$ (a) $P_{\text{Haar}}=0$, (b) $P_{\text{Haar}}=0.3$, (c) $P_{\text{Haar}}=0.7$, and (d) $P_{\text{Haar}}=1$.
I.3 Dynamics of CV with U(1)-symmetric(asymmetric) initial states at different $P_{\text{Haar}}$
As shown in Fig. S5, the circuit-averaged charge variance $\mathbb{E}[\sigma^{2}_{Q}]$ for all U(1)-symmetric initial states gradually increases until it reaches a plateau. This behavior contrasts with the overshooting dynamics observed in the entanglement asymmetry. Besides, the early-time dynamics of charge variance from different asymmetric initial states show no evidence of QME, as confirmed in Fig. S6. Two observations can be drawn from Figs. S5 and S6: (1) The time required to reach the plateau decreases as the density of random Haar gates increases. Furthermore, the charge variance saturates more quickly for the antiferromagnetic state compared to the ferromagnetic state. (2) The charge variance saturates to the same value for all U(1)-symmetric(asymmetric) initial states. This arises because the late-time density matrix for a small subsystem approaches an identity matrix. Here, $\mathbb{E}[\sigma^{2}_{Q}]$ is computed by averaging the charge variance, $\sigma^{2}_{Q}$, over $5000$ circuit configurations.
Figure S5: The dynamics of circuit-averaged CV, $\mathbb{E}[\sigma^{2}_{Q}]$, for different U(1)-symmetric initial states are examined for various values of $P_{\text{Haar}}$ with $L=16$. From left to right: the initial states are ferromagnetic, domain wall and antiferromagnetic states, respectively.
Figure S6: Dynamics of circuit averaged CV, $\mathbb{E}[\sigma^{2}_{Q}]$, for different U(1)-asymmetric initial states in a system with $L=16$. The top, middle, and bottom rows correspond to tilted ferromagnetic, tilted domain wall, and tilted antiferromagnetic initial states, respectively. The left and columns show results for $P_{\text{Haar}}=0.05$ and $P_{\text{Haar}}=0.1$, respectively.
I.4 The early and late time behavior of EA and CV in U(1) non-symmetric quantum circuits
In Table. 2, we provide a summary of the behavior of entanglement asymmetry and charge variance at early and late times for U(1)-asymmetric states in the circuit model. We observe that the quantum Mpemba effect occurs only in the early-time dynamics of entanglement asymmetry for tilted ferromagnetic and tilted domain wall states. Moreover, the entanglement asymmetry and charge variance approach distinct values, which are universal across all initial states, at late times.
Ferromagnetic Domain Wall Antiferromagnetic
$EA\ (early\ time)$ crossing crossing no crossing
$CV\ (early\ time)$ no crossing no crossing no crossing
$EA\ (late\ time)$ 0 0 0
$CV\ (late\ time)$ C C C
Table 2: The early- and late-time behavior of entanglement asymmetry (EA) and charge variance (CV) under the evolution of random circuit with $10\leq P_{\text{Haar}}<$ 1. The crossing in the behavior indicates the occurrence of the quantum Mpemba effect. The constant C can be evaluated as $\mathrm{tr}(\rho Q^{2})-\mathrm{tr}(\rho Q)^{2}$, where $\rho=\frac{I}{2^{L}}$ is the late-time density matrix for the subsystem.
II More numerical results of entanglement asymmetry for U(1) non-symmetric Hamiltonians
II.1 Dynamics of entanglement asymmetry for various U(1)-symmetric initial states
In Fig. S7, we show the dynamics of EA across all initial states (ferromagnetic, domain wall, and antiferromagnetic) under the evolution of $H_{2}$. The entanglement asymmetry shows evident overshooting at early times, and the U(1) symmetry remains broken in the subsystem at long times in all cases, consistent with the results described in the main text. As shown in Fig. S8, the peak of entanglement asymmetry also correlates with the strength of symmetry breaking, $1-\gamma$, for various initial states in the case of a non-integrable Hamiltonian. EA for ground state of $H_{2}$ is also shown for comparison. More importantly, the peak value of entanglement asymmetry is significantly higher than its late-time value.
Figure S7: EA as a function of time for various initial states and values of $\gamma$, with $L=12$. Panels (a) are based on the $H_{1}$, while panels (b), (c), and (d) correspond to $H_{2}$. Insets zoom in on the peak of EA for different values of $\gamma$ at early times, listed from bottom to top as: $\gamma=0.8,0.7,0.6,0.5,0.4,0.3,0.2,0.1$. (a),(c): Domain Wall state. (b): Ferromagnetic state. (d): Antiferromagnetic state.
Figure S8: Panels (a) and (b) show the peak value of EA, $(\Delta S_{L/3})_{max}$, and the ratio of the late-time EA, $\Delta S_{L/3}^{\infty}$, to $(\Delta S_{L/3})_{max}$ as a function of $1-\gamma$ for different initial states under $H_{2}$. F: Ferromagnetic state; DW: Domain Wall state; AF: Antiferromagnetic state; GS: Ground State.
II.2 Dynamics of entanglement entropy
To better explain the behavior of entanglement asymmetry, we examine the time evolution of entanglement entropy and explore its connection to entanglement asymmetry in this section. We present both $S_{L/3}$ and $S_{L/3,Q}$ as functions of time, with their difference yielding EA, for various initial states and Hamiltonians. As shown in Fig. S9, we observe that the entanglement entropy grows linearly with time before eventually reaching a plateau. The saturation time of the entanglement entropy is always greater than the time at which the entanglement asymmetry reaches its peak. In most cases, the peak of entanglement asymmetry occurs after $S_{L/3,Q}$ reaches its first peak. This observation is reflected in Fig. S9.
Figure S9: Dynamics of each individual term in the expression for the entanglement asymmetry $\Delta S_{L/3}$ with $L=12$ is shown. Solid lines represent $S_{L/3,Q}$, while dash lines correspond to the entanglement entropy $S_{L/3}$. From left to right: initial states are Ferromagnetic, Domain wall and Antiferromagnetic states, respectively. Panels (a), (b), and (c) are obtained using $H_{1}$, while panels (d), (e), and (f) are based on $H_{2}$.
II.3 Quantum Mpemba effect in U(1) non-symmetric Hamiltonian
When $\gamma=1$, the emergence of QME is observed in all cases, as shown in Figs. S10 and S11. Analogous to the discussion in the main text, we find the absence of QME in the non-integrable Hamiltonian across all initial states when $\gamma$ is tuned to above some thresholds.
Figure S10: EA dynamics at different values of $\gamma$ with $L=12$ are examined for tilted domain wall states. The calculation is based on $H_{1}$. Blue corresponds to $\theta=0.2\pi$, and red corresponds to $\theta=0.5\pi$. The increasing intensity of color reflects a stronger effect of symmetry breaking.
Figure S11: The dynamics of entanglement asymmetry are examined for different values of $\gamma$ with $L=12$ across three initial states: (a) tilted ferromagnetic states, (b) tilted domain wall states, and (c) tilted antiferromagnetic states. Panels (a)–(c) are based on the evolution under $H_{2}$. The blue curves correspond to $\theta=0.2\pi$, while the red curves correspond to $\theta=0.5\pi$. The increasing intensity of the color reflects the stronger effect of symmetry breaking.
II.4 Phase diagrams in U(1) non-symmetric Hamiltonians
As shown in Figs. S12 and S13, the phase diagrams for ferromagnetic and antiferromagnetic states obtained through $H_{2}$ are almost identical to those from $H_{1}$. A slight difference is observed in the case of the domain wall state. Since the phase boundaries are different for different initial states, a natural question arises – is the phase boundary only different in terms of $\theta$ or also different when translating $\theta$ to initial EA. In Fig. S14, we replace $\theta$ with the initial value of the entanglement asymmetry, $\Delta S_{L/4}^{ini}$, on the $y$-axis. We find that the phase boundary is still different in terms of initial EA for different types of initial states.
Figure S12: A schematic phase diagram illustrating the dependence of EA dynamical patterns on the parameters $\theta$ and $1-\gamma$ for a domain wall state under the evolution of $H_{1}$. Those black dots are obtained through numerical simulation.
Figure S13: Panels (a)-(c) show schematic phase diagrams depicting the dependence of EA dynamical patterns on $\theta$ and $1-\gamma$ for different initial states. All panels are based on $H_{2}$. (a) Ferromagnetic state, (b) Domain Wall state, (c) Antiferromagnetic state. Those black dots are obtained through numerical simulation.
Figure S14: Panels (a)-(f) show schematic phase diagrams depicting the dependence of EA dynamical patterns on the initial value of EA, $\Delta S_{L/4}^{ini}$ and $1-\gamma$ for different initial states and Hamiltonians. From left to right: initial states are Ferromagnetic, Domain wall and Antiferromagnetic states, respectively. Panels (a), (b), and (c) are obtained using $H_{1}$, while panels (d), (e), and (f) are based on $H_{2}$.
II.5 Late-time entanglement asymmetry
The late-time entanglement asymmetry is obtained by averaging $\Delta S_{L/4}$ over $2000$ time points between $t_{1}=2000$ and $t_{2}=40000$, as described in the main text. In S15, we compute the late-time EA for different initial states under the evolution of $H_{1}$ and $H_{2}$. We find two general trends: (1) The late-time value of EA increases with the strength of symmetry breaking $1-\gamma$. (2) For states with higher initial asymmetry, the entanglement asymmetry at late times is greater and increases more rapidly compared to states with lower initial asymmetry.
Figure S15: The late-time entanglement asymmetry, $\Delta S_{L/4}^{\infty}$, as a function of $1-\gamma$ with $L=12$. From left to right: the initial states are tilted ferromagnetic, tilted domain wall and tilted antiferromagnetic states, respectively. Panels (a), (b), and (c) correspond to results obtained using $H_{1}$, while panels (d), (e), and (f) are based on $H_{2}$.
III More numerical results of charge variance for U(1) non-symmetric Hamiltonians
Figure S16: The time evolution of the expectation value of the total spin in the $z$-direction, $\langle Q_{tot}\rangle$, is investigated for various values of $\gamma$ with a ferromagnetic initial state, where $L=12$. Here, $\hat{Q}_{tot}=\sum_{i=1}^{L}\sigma_{i}^{z}$. Left: $H_{1}$. Right: $H_{2}$.
Figure S17: The variance of $Q_{tot}$, $\sigma_{Q}^{2}$, as a function of time for different values of $\gamma$ with $L=12$. Top row: $H_{1}$. Bottom row: $H_{2}$. From left to right: Ferromagnetic, Domain Wall, and Antiferromagnetic states.
III.1 Expectation values of $\hat{Q}$ and $\sigma_{Q}^{2}$
The expectation value of the total spin in the $z$-direction as a function of time is shown in Fig. S16 for ferromagnetic states. At the start, $Q_{tot}=L=12$ since all the spins are aligned upward. Over time, the state progressively loses memory of its initial configuration, and $Q_{tot}$ settles into oscillations around 00. For domain wall and antiferromagnetic states, $\langle\hat{Q}_{tot}(t)\rangle=0$ is a constant due to symmetry argument which are elaborated below.
Suppose we have a transformation $T$:
$T:$ $\quad\sigma^{z}\rightarrow-\sigma^{z},\quad\sigma^{y}\rightarrow-\sigma^{y}$ (S1)
It is straightforward to verify that the Hamiltonians remain invariant under this transformation $T$. Thus, we have
$\langle\hat{Q}_{\text{tot}}(t)\rangle$ $=\langle\psi(t)|\hat{Q}_{\text{tot}}|\psi(t)\rangle$
$=\langle\psi(0)|e^{iHt}\hat{Q}_{\text{tot}}e^{-iHt}|\psi(0)\rangle$
$=\langle\psi(0)|T^{\dagger}Te^{iHt}T^{\dagger}T\hat{Q}_{\text{tot}}T^{\dagger}Te^{-iHt}T^{\dagger}T|\psi(0)\rangle$
$=\langle\psi(0)|T^{\dagger}e^{iHt}(-\hat{Q}_{\text{tot}})e^{-iHt}T|\psi(0)\rangle$ (S2)
$=\langle\psi_{p}(0)|e^{iHt}(-\hat{Q}_{\text{tot}})e^{-iHt}|\psi_{p}(0)\rangle$
$=-\langle\psi_{p}(t)|\hat{Q}_{\text{tot}}|\psi_{p}(t)\rangle$
$=-\langle\psi(t)|\hat{Q}_{\text{tot}}|\psi(t)\rangle$
$=0$
Here $|\psi_{p}(t)\rangle$ is related to $|\psi(t)\rangle$ by a permutation of site indices. Throughout the derivation, we use the fact that $T\hat{Q}_{\text{tot}}T^{\dagger}=-\hat{Q}_{\text{tot}}$. Therefore, $\langle\hat{Q}_{tot}(t)\rangle$ is strictly zero for domain wall and antiferromagnetic initial states.
Fig. S17 shows the time dependence of the variance of total charge operator for different initial states under $H_{1}$ and $H_{2}$. Notably, the variance of the charge typically saturates later than the entanglement asymmetry. The charge variance also characterizes the U(1) symmetry breaking in the evolved state from some aspects, similar as EA explored in the main text. However, the symmetry breaking dynamical behaviors show distinct patterns in the two metrics. EA shows an evident overshooting while the charge variance directly saturates. The differences may provide further insight into the multifaceted physics of symmetry breaking.
III.2 Dynamics of charge variance for different U(1)-asymmetric initial states
Figures S18 and S19 show that, for charge variance evolved from the tilted domain wall and tilted antiferromagnetic states, the charge variance is consistently larger for more asymmetric states (large $\theta$) than for less asymmetric ones (small $\theta$). However, this trend does not apply to the initial tilted ferromagnetic state, where the monotonic relationship of charge variance with respect to $\theta$ is reversed at early times. This results in an early-time crossing, which becomes more pronounced as $1-\gamma$ increases. These behaviors are observed in both integrable and non-integrable Hamiltonians, highlighting their universality across different dynamical regimes.
Figure S18: Temporal evolution of the charge variance, $\sigma_{Q}^{2}$, under the integrable Hamiltonian $H_{1}$ for a system size $L=12$. The top row corresponds to tilted ferromagnetic states, the middle row to tilted domain wall states, and the bottom row to tilted antiferromagnetic states. From left to right, the columns show results for $\gamma=0.9$, $0.8$, and $0.7$.
Figure S19: Time evolution of the charge variance, $\sigma_{Q}^{2}$, under the non-integrable Hamiltonian $H_{2}$ for a system size $L=12$. The top row corresponds to tilted ferromagnetic states, the middle row to tilted domain wall states, and the bottom row to tilted antiferromagnetic states. From left to right, the columns show results for $\gamma=0.9$, $0.8$, and $0.7$.
III.3 Dynamics of charge sector probability distributions
Figure S20 displays the probability distribution $P_{Q}$ across various charge sectors $Q$ for the evolved state $|\psi(t)\rangle$ at different times for initial tilted ferromagnetic states. Here, $P_{Q}$ at time $t$ is defined as $\sum_{q=Q}|\langle\psi(t)|\psi_{q}\rangle|^{2}$, where $|\psi_{q}\rangle$ represents the basis wave function corresponding to charge $q$. For more asymmetric initial states (larger $\theta$), the charge distribution shrinks, leading to a decrease in charge variance. In contrast, for less asymmetric states, the distribution spreads across a broader range of charge sectors, increasing the charge variance. This behavior accounts for the crossing phenomenon observed in Figs. S18 and S19.
Figure S20: Time evolution of the probability distribution, $P_{Q}$, for each charge sector $Q$ under the Hamiltonian $H_{1}$ with $\gamma=0.6$. The charge sectors range from $Q=-12$ to $Q=12$ in increments of $2$. The rows, from top to bottom, correspond to time points $t=0$, $0.75$, $1.5$, $3$, and $4$. The columns, from left to right, represent different tilted ferromagnetic states with $\theta=0.2\pi$, $0.3\pi$, $0.4\pi$, and $0.5\pi$.
III.4 Phase diagrams in U(1) non-symmetric Hamiltonian
Figure S21 presents schematic 2-dimensional phase diagrams that capture the dependence of early-time CV dynamics on the parameters $\theta$ and $1-\gamma$ for the ferromagnetic state evolving under $H_{1}$ and $H_{2}$. The diagram derived from $H_{2}$ closely resembles those obtained from $H_{1}$. The red region indicates where CV initially grows, While the blue region corresponds to an initial decrease in CV. At $\gamma=1$ (left boundary), the charge variance remains constant and equal to its initial value.
Figure S21: Panel (a) and (b) present schematic phase diagrams depicting the dependence of charge variance on $\theta$ and $1-\gamma$ with ferromagnetic state evolving under $H_{1}$ (left) and $H_{2}$ (right).
III.5 Late-time charge variance
The late-time charge variance, $\sigma_{Q}^{2}(t\rightarrow\infty)$, is computed in a manner similar to the late-time entanglement asymmetry. We examine the late-time charge variance for different initial states under the evolution of both $H_{1}$ and $H_{2}$, as illustrated in Fig. S22. It is observed that the late-time charge variance increases with the tilted angle $\theta$ for domain wall and antiferromagnetic states at a fixed $\gamma$. In contrast, for ferromagnetic states, the late-time charge variance decreases as $\theta$ increases, which is highly non-trivial. This behavior, also observed at early times (as seen in the crossing in Figs. S18 and S19), continues into the late-time regime.
Figure S22: The late-time charge variance, $\sigma_{Q\infty}^{2}$, as a function of $1-\gamma$ with $L=12$. From left to right: the initial states are tilted ferromagnetic, tilted domain wall and tilted antiferromagnetic states, respectively. Panels (a), (b), and (c) correspond to results obtained using $H_{1}$, while panels (d), (e), and (f) are based on $H_{2}$.
IV Analytical results of charge variance at early times
The early-time dynamics of charge variance, $\sigma_{Q}^{2}=\langle Q^{2}\rangle-\langle Q\rangle^{2}$, can be expanded as:
$\sigma_{Q}^{2}(t)\approx\sigma_{Q}^{2}(0)+t\left.\frac{d\sigma_{Q}^{2}}{dt}\right|_{0}+\frac{t^{2}}{2}\left.\frac{d^{2}\sigma_{Q}^{2}}{dt^{2}}\right|_{0}+...$ (S3)
where the ellipsis represents higher order terms that are neglected. Here, $\left.\frac{d\sigma_{Q}^{2}}{dt}\right|_{0}$ and $\left.\frac{d^{2}\sigma_{Q}^{2}}{dt^{2}}\right|_{0}$ denote the first and second derivatives of the charge variance evaluated at $t=0$, respectively. The initial state is chosen to be a tilted ferromagnetic state. Using the Heisenberg equation of motion, the first derivative of the charge variance at $t=0$ is given by:
$\left.\frac{d\sigma_{Q}^{2}}{dt}\right|_{0}=i\langle[H,Q^{2}]\rangle_{0}-2i\langle Q\rangle_{0}\langle[H,Q]\rangle_{0}$ (S4)
To evaluate $[H,Q]$, where the Hamiltonian is:
$H=-\frac{1}{4}\sum_{i=1}^{L}(\sigma_{i}^{x}\sigma_{i+1}^{x}+\gamma\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta_{1}\sigma_{i}^{z}\sigma_{i+1}^{z})$ (S5)
and the charge operator is $Q=\sum_{j=1}^{L}\sigma_{j}^{z}$, We compute:
$[H,Q]$ $=$ $[-\frac{1}{4}\sum_{i=1}^{L}(\sigma_{i}^{x}\sigma_{i+1}^{x}+\gamma\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta_{1}\sigma_{i}^{z}\sigma_{i+1}^{z}),\sum_{j=1}^{L}\sigma_{j}^{z}]$ (S6)
$=$ $-\frac{1}{4}\sum_{j}[\sigma_{j-1}^{x}\sigma_{j}^{x}+\sigma_{j}^{x}\sigma_{j+1}^{x}+\gamma\sigma_{j-1}^{y}\sigma_{j}^{y}+\gamma\sigma_{j}^{y}\sigma_{j+1}^{y},\sigma_{j}^{z}]$ (S7)
Simplifying, we obtain:
$[H,Q]=\frac{i}{2}(1-\gamma)\sum_{j}(\sigma_{j}^{x}\sigma_{j+1}^{y}+\sigma_{j}^{y}\sigma_{j+1}^{x})$ (S8)
Since $\langle\sigma_{j}^{y}\rangle_{0}=0$, it follows that $\langle[H,Q]\rangle_{0}=0$. Similarly, one can show that $\langle[H,Q^{2}]\rangle_{0}=0$. Consequently, the linear term in the expansion Eq. S3 vanishes. The quadratic term in the expansion is determined by the second derivative of the charge variance:
$\left.\frac{d^{2}\sigma_{Q}^{2}}{dt^{2}}\right|_{0}=-\langle[H,[H,Q^{2}]]\rangle_{0}+2\langle Q\rangle_{0}\langle[H,[H,Q]]\rangle_{0}$ (S9)
To compute $[H,[H,Q]]$, we evaluate:
$[H,[H,Q]]$ $=$ $[-\frac{1}{4}\sum_{i=1}^{L}(\sigma_{i}^{x}\sigma_{i+1}^{x}+\gamma\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta_{1}\sigma_{i}^{z}\sigma_{i+1}^{z}),\frac{i}{2}(1-\gamma)\sum_{j}(\sigma_{j}^{x}\sigma_{j+1}^{y}+\sigma_{j}^{y}\sigma_{j+1}^{x})]$ (S10)
$=$ $\sum_{j}\frac{1}{2}(1-\gamma)((1-\gamma)\sigma_{j}^{z}+\sigma_{j-1}^{x}\sigma_{j}^{z}\sigma_{j+1}^{x}-\gamma\sigma_{j-1}^{y}\sigma_{j}^{z}\sigma_{j+1}^{y})$ (S11)
-- $\sum_{j}\frac{1}{4}\Delta_{1}(1-\gamma)(\sigma_{j-1}^{x}\sigma_{j}^{x}\sigma_{j+1}^{z}+\sigma_{j-1}^{z}\sigma_{j}^{x}\sigma_{j+1}^{x})$ (S12)
Using $\langle\sigma_{j}^{x}\rangle_{0}=\sin{\theta}$ ) and $\langle\sigma_{j}^{z}\rangle_{0}=\cos{\theta}$ ). We find:
$\langle[H,[H,Q]]\rangle_{0}=\frac{1}{2}(1-\gamma)((1-\gamma)L\cos{\theta}+L\cos{\theta}\sin^{2}{\theta})-\frac{1}{2}\Delta_{1}(1-\gamma)L\cos{\theta}\sin^{2}{\theta}$ (S13)
Next, we evaluate the term $[H,[H,Q^{2}]]$. Expanding this expression, we obtain:
$[H,[H,Q^{2}]]$ $=$ $[H,Q[H,Q]]+[H,[H,Q]Q]$ (S14)
$=$ $Q[H,[H,Q]]+[H,[H,Q]]Q+2[H,Q]^{2}$ (S15)
The expectation value of the third term is:
$\langle[H,Q]^{2}\rangle_{0}$ $=$ $-\frac{1}{4}(1-\gamma)^{2}\langle\sum_{i}(\sigma_{i}^{x}\sigma_{i+1}^{y}+\sigma_{i}^{y}\sigma_{i+1}^{x})\sum_{j}(\sigma_{j}^{x}\sigma_{j+1}^{y}+\sigma_{j}^{y}\sigma_{j+1}^{x})\rangle_{0}$ (S16)
$=$ $-\frac{1}{2}(1-\gamma)^{2}L(1+\sin^{2}{\theta})$ (S17)
The first term, $\langle Q[H,[H,Q]]\rangle_{0}$, is computed as:
$\langle Q[H,[H,Q]]\rangle_{0}$ $=$ $\langle\frac{1}{2}(1-\gamma)^{2}\sum_{i,j}\sigma_{i}^{z}\sigma_{j}^{z}+\frac{1}{2}(1-\gamma)\sum_{i,j}\sigma_{i}^{z}\sigma_{j}^{z}(\sigma_{j-1}^{x}\sigma_{j+1}^{x}-\gamma\sigma_{j-1}^{y}\sigma_{j+1}^{y})\rangle_{0}$ (S18)
-- $\langle\sum_{i,j}\frac{1}{4}\Delta_{1}(1-\gamma)(\sigma_{i}^{z}\sigma_{j-1}^{x}\sigma_{j}^{x}\sigma_{j+1}^{z}+\sigma_{i}^{z}\sigma_{j-1}^{z}\sigma_{j}^{x}\sigma_{j+1}^{x})\rangle_{0}$ (S19)
$=$ $\frac{1}{2}(1-\gamma)^{2}(L+L(L-1)\cos^{2}{\theta})+\frac{1}{2}(1-\gamma)L(L-3)\sin^{2}{\theta}\cos^{2}{\theta}+\frac{1}{2}(1-\gamma)L\sin^{2}{\theta}$ (S20)
-- $\frac{1}{2}\Delta_{1}(1-\gamma)(L(L-3)\sin^{2}{\theta}\cos^{2}{\theta}+L\sin^{2}{\theta})$ (S21)
Here, we have used $\langle Q^{2}\rangle_{0}=L+L(L-1)\cos^{2}{\theta}$. The second term in Eq. S18 has two contributions from $i=j$ and $i\neq j$, $j-1$, and $j+1$, while the third term vanishes due to $\langle\sigma_{j}^{y}\rangle_{0}=0$. Moreover, $\langle Q[H,[H,Q]]\rangle_{0}=\langle[H,[H,Q]]Q\rangle_{0}$. Collecting all terms in Eq. S15, the first term in Eq. S9 becomes
$-\langle[H,[H,Q^{2}]]\rangle_{0}$ $=$ $-(1-\gamma)^{2}(L+L(L-1)\cos^{2}{\theta})-(1-\gamma)L(L-3)\sin^{2}{\theta}\cos^{2}{\theta}-(1-\gamma)L\sin^{2}{\theta}$ (S22)
$+$ $\Delta_{1}(1-\gamma)(L(L-3)\sin^{2}{\theta}\cos^{2}{\theta}+L\sin^{2}{\theta})+(1-\gamma)^{2}L(1+\sin^{2}{\theta})$ (S23)
$=$ $-(1-\gamma)^{2}(L^{2}\cos^{2}{\theta}-L)-(1-\gamma)(L^{2}-3L)\sin^{2}{\theta}\cos^{2}{\theta}-(1-\gamma)L\sin^{2}{\theta}$ (S24)
$+$ $\Delta_{1}(1-\gamma)(L^{2}-3L)\sin^{2}{\theta}\cos^{2}{\theta}+\Delta_{1}(1-\gamma)L\sin^{2}{\theta}$ (S25)
The second term in Eq. S9 is computed as:
[…]
$=$ $L^{2}(1-\gamma)^{2}\cos^{2}{\theta}+L^{2}(1-\gamma)\cos^{2}{\theta}\sin^{2}{\theta}-\Delta_{1}(1-\gamma)L^{2}\cos^{2}{\theta}\sin^{2}{\theta}$ (S27)
Combining all terms, the second derivative of the charge variance is:
$\left.\frac{d^{2}\sigma_{Q}^{2}}{dt^{2}}\right|_{0}$ $=$ $-\langle[H,[H,Q^{2}]]\rangle_{0}+2\langle Q\rangle_{0}\langle[H,[H,Q]]\rangle_{0}$ (S28)
$=$ $-(1-\gamma)^{2}(L^{2}\cos^{2}{\theta}-L))-(1-\gamma)(L^{2}-3L)\sin^{2}{\theta}\cos^{2}{\theta}-(1-\gamma)L\sin^{2}{\theta}$ (S29)
$+$ $\Delta_{1}(1-\gamma)(L^{2}-3L)\sin^{2}{\theta}\cos^{2}{\theta}+\Delta_{1}(1-\gamma)L\sin^{2}{\theta}$ (S30)
$+$ $L^{2}(1-\gamma)^{2}\cos^{2}{\theta}+L^{2}(1-\gamma)\cos^{2}{\theta}\sin^{2}{\theta}-\Delta_{1}(1-\gamma)L^{2}\cos^{2}{\theta}\sin^{2}{\theta}$ (S31)
$=$ $L(1-\gamma)^{2}+3L(1-\gamma)\sin^{2}{\theta}\cos^{2}{\theta}-L(1-\gamma)\sin^{2}{\theta}$ (S32)
-- $3\Delta_{1}L(1-\gamma)\sin^{2}{\theta}\cos^{2}{\theta}+\Delta_{1}L(1-\gamma)\sin^{2}{\theta}$ (S33)
The early-time growth of the charge variance for different values of $\theta$ is then approximated by:
$\sigma_{Q}^{2}(t)/L=\sin^{2}{\theta}+\frac{t^{2}}{2}(1-\gamma)\left(1-\gamma+(1-\Delta_{1})(3\sin^{2}\theta\cos^{2}\theta-\sin^{2}\theta\right))$ (S34)
Here, the initial charge variance, $\sigma_{Q}^{2}(0)$, is replaced with $L(1-\cos^{2}{\theta})$. We can now compare this analytical result with numerical simulations at early times. We find that the charge variance obtained from both methods is in good agreement at early times, as confirmed in Fig. S23.
Figure S23: Comparison of the early-time behavior of the charge variance, $\sigma_{Q}^{2}$, between analytical solutions (dashed lines) and numerical simulations (solid lines). Results are obtained from tilted ferromagnetic states under the Hamiltonian $H_{1}$ with $L=12$. Panel (a): $\gamma=0.7$. Panel (b): $\gamma=0.6$.

### v3

@@
 Entanglement asymmetry has emerged as a powerful tool for characterizing symmetry breaking in quantum many-body systems.
-In this Letter, we explore how symmetry is dynamically broken through the lens of entanglement asymmetry in two distinct scenarios: a non-symmetric random quantum circuit and a non-symmetric Hamiltonian quench, with a particular focus on U(1) symmetry.
-In the former case, the symmetry is initially broken and subsequently restored, whereas in the latter case, symmetry remains broken in the subsystem at late times, consistent with the principles of quantum thermalization.
+In this Letter, we explore how symmetry is dynamically broken through the lens of entanglement asymmetry in two distinct scenarios: a non-symmetric Hamiltonian quench and a non-symmetric random quantum circuit, with a particular focus on U(1) symmetry.
+In the former case, symmetry remains broken in the subsystem at late times, whereas in the latter case, the symmetry is initially broken and subsequently restored, consistent with the principles of quantum thermalization.
 Notably, the growth of entanglement asymmetry exhibits unexpected overshooting behavior at early times in both contexts, contrasting with the behavior of charge variance.
@@
 Symmetry properties are also closely related to the concept of quantum thermalization [2, 3, 4, 5, 6] for generic quantum many-body systems.
-In general, when a closed quantum system evolves with a chaotic Hamiltonian, the reduced density matrix of a small subsystem $a$ thermalizes to the equilibrium finite-temperature state: $\rho_{a}\propto e^{-\beta\hat{H}_{a}}$ where $\hat{H}_{a}$ is the Hamiltonian of the subsystem.
-Symmetry is restored at later times for symmetric Hamiltonian $\hat{H}_{a}$, since $[\hat{Q}_{a},\rho_{a}]=0$ where $\hat{Q}_{a}$ represents the corresponding symmetry generator.
+In general, when a closed quantum system evolves with a chaotic Hamiltonian, the reduced density matrix of a small subsystem $a$ thermalizes to the equilibrium finite-temperature state: $\rho_{a}\propto e^{-\beta\hat{H}_{a}}$ where $\hat{H}_{a}$ is the Hamiltonian of the subsystem and $\beta$ is a Lagrangian multiplier determined by the initial energy density.
+Symmetry is restored at later times for symmetric Hamiltonian $\hat{H}_{a}$, since $\hat{Q}_{a},\rho_{a}]=0$ where $\hat{Q}_{a}$ represents the corresponding symmetry generator.
 However, if $\hat{H}_{a}$ does not respect the symmetry, the reduced density matrix $\rho_{a}$ at late times is non-commuting with $\hat{Q}_{a}$.
 In this case, symmetry breaking persists even if the system begins in a symmetric state.
-Despite the richness of the late-time behavior, early-time dynamics have also garnered significant attention in recent years, particularly due to the novel Mpemba effect [7], which demonstrates that hot water freezes faster than cold water.
-Both classical and quantum versions of the Mpemba effect have been widely explored in various systems [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
-Recently, quantum Mpemba effect (QME) is reported in quantum integrable systems and chaotic systems [24, 25, 26].
-Notably, U(1)-symmetry restoration occurs more rapidly for more asymmetric initial states under the U(1)-symmetric Hamiltonian quench [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37].
+It is important to distinguish this explicit breaking from spontaneous symmetry breaking (SSB).
+In the thermodynamic limit, SSB enables a low-temperature symmetric initial state to evolve dynamically into a symmetry-broken steady state.
+In contrast, for finite-size systems, SSB cannot occur, and the equilibrium state must retain the symmetry of the Hamiltonian.
+Apart from the richness of the late-time behavior, early-time dynamics have also garnered significant attention.
+For example, the Mpemba effect [7], which claims that hot water freezes faster than cold water, has been widely explored in both classical and quantum contexts [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
+Recently, quantum Mpemba effect (QME) is reported in integrable systems and chaotic systems [24, 25, 26].
+QME refers to the phenomenon where, during the relaxation toward a steady state value, the time evolution curves of a physical quantity for different initial conditions cross each other.
+For instance, U(1)-symmetry restoration occurs more rapidly for more asymmetric initial states under the U(1)-symmetric Hamiltonian quench [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37].
 This finding was subsequently explored in various other settings [38, 39, 40, 41, 42, 43, 44, 45, 46, 47] and experimentally realized on a trapped-ion quantum simulator [48].
@@
 Addressing these questions offers a more comprehensive understanding of symmetry and symmetry breaking in quantum many-body systems.
-Figure 1: (a) Schematic illustration of a non-symmetric random circuit with 6 qubits.
-Gates are arranged in the even-odd brick-wall pattern.
-The blue and red rectangles represent U(1)-symmetric and random Haar gates, respectively.
-The basis for the U(1)-symmetric gate is listed in the following order: $|00\rangle$, $|01\rangle$, $|10\rangle$ and $|11\rangle$.
-(b) The circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]$, as a function of time with the antiferromagnetic initial state at different values of $P_{\text{Haar}}$.
-(c) The peak value, $\mathbb{E}[\Delta S_{L/4}]_{max}$, as a function of $P_{\text{Haar}}$.
-All three curves follow a power law $y=ax^{b}$.
-F: Ferromagnetic state ($a=1.4$, $b=0.4$); DW: Domain Wall state ($a=2.7$, $b=0.8$); AF: Antiferromagnetic state ($a=1.9$, $b=0.9$).
 In this Letter, we investigate and compare the dynamics of symmetry breaking with two distinct models: a non-symmetric random circuit [50] and a non-symmetric Hamiltonian evolution, each with different symmetric and asymmetric initial states.
@@
 Consequently, $\rho_{a,Q}$ is block diagonal in the eigenbasis of $\hat{Q}_{a}$.
-The EA satisfies two key properties: (1) $\Delta S_{a}\geq 0$ since the EA is defined as the relative entropy between $\rho_{a,Q}$ and $\rho_{a}$.
+The EA satisfies two key properties: (1) $\Delta S_{a}\geq 0$ since the EA is the relative entropy between $\rho_{a,Q}$ and $\rho_{a}$.
 (2) $\Delta S_{a}=0$ if and only if $\rho_{a,Q}=\rho_{a}$.
@@
 In parallel with the analysis of EA, we also compute the charge variance (CV) $\sigma^{2}_{Q}=\langle\hat{Q}^{2}\rangle-\langle\hat{Q}\rangle^{2}$, where $\hat{Q}=\sum_{i=1}^{L}\sigma_{i}^{z}$.
-This quantity serves as a measure of charge fluctuations within the system, offering a complementary perspective on the dynamics of symmetry breaking.
-In the case of non-symmetric random circuits, we show that U(1) symmetry for a small subsystem can still be restored regardless of whether the initial state is U(1)-symmetric or asymmetric.
-As a result, EA exhibits overshooting at early times, i.e.
-there is a peak of EA at an early time that is much larger than the saturating EA value at late times.
-Additionally, QME appears at early times unless all U(1)-symmetric gates are replaced by random Haar gates, where EA dynamics are exactly the same for different U(1)-asymmetric initial states.
-For a non-symmetric Hamiltonian evolution, we find that U(1) symmetry can not be restored in a subsystem.
-This behavior can be explained by the late-time reduced density matrix relaxing to the form $e^{-\beta\hat{H}_{a}}$, where $\hat{H}_{a}$ explicitly includes symmetry-breaking terms.
-In this scenario, the EA still shows nontrivial overshooting at early times, contrasting to other symmetry-breaking measures such as charge variance.
-Furthermore, early-time EA dynamics from asymmetric initial states display distinct behaviors due to the competition between two symmetry-breaking contributions related to the initial state and the system Hamiltonian.
-Moreover, the QME originated from symmetric evolution disappears when the strength of symmetry breaking in the evolution exceeds some thresholds.
-Setup.— To characterize symmetry breaking in these systems, we consider three initial states: the ferromagnetic state $|000...0\rangle$, the antiferromagnetic state $|0101..1\rangle$ and the domain-wall state $|000..111\rangle$, where the domain wall is positioned at the center of the chain.
-To incorporate the effect of symmetry breaking in the initial state, we introduce tilted ferromagnetic states, as discussed in earlier studies [24, 25].
-The state is defined as
-$|\psi_{i}(\theta)\rangle=e^{-i\frac{\theta}{2}\sum_{j}\sigma_{j}^{y}}|000...0\rangle$ (2)
-where $\sigma_{j}^{y}$ is the Pauli matrix in $y$-direction acting on the $j$-th qubit, and $\theta$ is a tuning parameter that controls the strength of symmetry breaking in the initial state.
-When $\theta=0$, Eq.
-(2) is U(1)-symmetric, resulting in a vanishing EA.
-As $\theta$ increases, the EA grows, reaching its maximum value at $\theta=\pi/2$.
-The tilted antiferromagnetic and tilted domain wall states are constructed in a similar manner.
-The random circuit in Fig.
-1 (a) consists of two-qubit random U(1)-symmetric gates and random Haar gates, arranged in a brick-wall fashion.
-The exact matrix form of U(1)-symmetric gate is illustrated in Fig.
-1 (a), where each block is randomly drawn from the Haar measure [57, 58, 59].
-The effect of symmetry breaking depends on the density (doped probability) of random Haar gates without U(1) symmetry, denoted as $P_{\text{Haar}}$.
-The time unit in the circuit is defined by the application of two consecutive layers of gates.
-The initial state $|\psi_{i}(\theta)\rangle$ evolves under the random unitary dynamics and $\mathbb{E}[\Delta S_{a}]$ is computed by averaging $\Delta S_{a}$ over $5000$ circuit configurations.
-We also investigate Hamiltonian dynamics where the state $|\psi_{i}(\theta)\rangle$ undergoes unitary evolution given by $e^{-iHt}|\psi_{i}(\theta)\rangle$, and the Hamiltonian is defined as
-$H=$ $-\frac{1}{4}\sum_{j=1}^{L}\Big{[}\sigma_{j}^{x}\sigma_{j+1}^{x}+\gamma\sigma_{j}^{y}\sigma_{j+1}^{y}+\Delta_{1}\sigma_{j}^{z}\sigma_{j+1}^{z}\Big{]}$
-$-\Delta_{2}\sum_{j=1}^{L}\Big{[}\sigma_{j}^{x}\sigma_{j+2}^{x}+\sigma_{j}^{y}\sigma_{j+2}^{y}+\sigma_{j}^{z}\sigma_{j+2}^{z}\Big{]}.$
-Here, $\Delta_{1}$ and $\Delta_{2}$ are the coefficients for nearest-neighbor and next-nearest-neighbor interactions, respectively.
-$\Delta_{2}$ introduces non-integrability, and $\gamma$ controls the strength of symmetry breaking.
-Periodic boundary conditions are imposed in both contexts.
-Figure 2: The circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]$, as a function of time for different values of $P_{\text{Haar}}$.
-Blue: U(1) EA.
-Green: $Z_{2}$ EA.
-Panels (a)-(d) correspond to different values of $P_{\text{Haar}}$ (a) $P_{\text{Haar}}=0$, (b) $P_{\text{Haar}}=0.3$, (c) $P_{\text{Haar}}=0.7$, and (d) $P_{\text{Haar}}=1$, respectively.
-U(1)-Symmetric (Asymmetric) States with U(1) Non-Symmetric Random Circuit.— All numerical simulations are performed using the TensorCircuit-NG package [60].
-The circuit under investigation consists of 16 qubits.
-We evaluate the EA at different doping probabilities of random Haar gates, using an antiferromagnetic initial state.
-We observe that, at later times, all EAs approach zero, as illustrated in Fig.
-1 (b).
-This behavior can be understood in the context of quantum thermalization and information scrambling [61, 62, 63, 64], where the reduced density matrix of the subsystem is a fully mixed state for the random circuit cases, as long as the subsystem size does not exceed half of the total system size.
-Additionally, all EAs display a peak at some early time steps.
-For all probabilities chosen in Fig.
-1 (b), all EAs reach their maximum after only a few layers of unitaries.
-The rate of symmetry restoration also depends on the initial state.
-In the supplementary material (SM), we find that symmetry restoration occurs more quickly for antiferromagnetic or domain wall states than for ferromagnetic states, due to the larger Hilbert space sector of the initial states in the former cases.
-In Fig.
-1 (c), we reveal that the peak of the circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]_{max}$ follows a power-law with respect to $P_{\text{Haar}}$.
-Next, we examine the dynamics from U(1)-asymmetric initial states, i.e.
-a tilted ferromagnetic state.
-We compute the EA for both U(1) symmetry with $\hat{Q_{a}}=\sum_{i\in a}\sigma_{i}^{z}$ and $Z_{2}$ symmetry with $\hat{Q_{a}}=\prod_{i\in a}\sigma_{i}^{z}$.
-As depicted in Fig.
-2 (a), for $P_{\text{Haar}}=0$, we clearly notice the emergence of QME in U(1) case.
-Surprisingly, we also find that the QME appears in the $Z_{2}$ probe, which does not contradict previous study [25] suggesting the absence of QME in $Z_{2}$-symmetric circuits.
-Even though U(1)-symmetric gates are also $Z_{2}$ symmetric, there is no off-diagonal coupling between $|00\rangle$ and $|11\rangle$, leading to different thermalization rates between two $Z_{2}$ charge sectors ($Q_{a}=\pm 1$), and thus resulting in QME.
+This quantity serves as a measure of charge fluctuations within the system, offering a complementary perspective on symmetry breaking.
 Figure 1: EA as a function of time with (a) ferromagnetic and (b) antiferromagnetic states for different values of $\gamma$ under $H_{1}$.
@@
 GS denotes the value of EA calculated from the ground state of $H_{1}$.
-As we replace a portion of U(1)-symmetric gates with random Haar gates, QME remains evident with a finite number of random Haar gates.
-However, when the circuit consists entirely of random Haar gates, all charge sectors thermalize at the same rate after circuit averaging, and QME disappears.
-U(1)-Symmetric Initial States with U(1) Non-Symmetric Hamiltonian.— Here, we investigate the dynamics of symmetry breaking under an integrable Hamiltonian $H_{1}$ with $\Delta_{1}=0.4$ and $\Delta_{2}=0$, and a non-integrable Hamiltonian $H_{2}$ with $\Delta_{1}=0.4$ and $\Delta_{2}=0.05$, with system size 12 sites.
-As revealed in Fig.
-3 (a) and (b), we calculate EA for various Hamiltonian symmetry-breaking values $\gamma$ and observe that EAs also exhibit peaks at early times that are much larger than steady values.
-Furthermore, the peak value of the EA, $(\Delta S_{L/3})_{max}$, is found to be correlated with the strength of symmetry breaking, $1-\gamma$, for different symmetric initial states as shown in Fig.
-3 (c) where EA of the ground state of $H_{1}$ follows the same trend.
-Notably, the peak heights nearly coincide between the ferromagnetic and domain wall states, as the early-time peak primarily depends on the local configurations of the initial state.
+For a non-symmetric Hamiltonian evolution, we find that U(1) symmetry can not be restored in a subsystem which can be explained by the late-time reduced density matrix relaxing to the form $e^{-\beta\hat{H}_{a}}$, where $\hat{H}_{a}$ explicitly includes symmetry-breaking terms.
+In this scenario, the EA shows nontrivial overshooting at early times, characterized by a peak in EA that significantly exceeds its late-time saturation value.
+This behavior contrasts with other symmetry-breaking measures like charge variance and mirrors the thermal overshooting of the classical Mpemba effect [11], where systems transiently exceed their equilibrium temperature.
+Furthermore, the QME originated from symmetric evolution disappears when the strength of symmetry breaking in the evolution exceeds some thresholds.
+In the case of non-symmetric random circuits, we show that U(1) symmetry for a small subsystem can still be restored regardless of the initial states.
+As a result, EA also exhibits overshooting at early times.
+Additionally, QME appears at early times unless all U(1)-symmetric gates are replaced by random Haar gates, where EA dynamics are the same for different U(1)-asymmetric initial states.
 Figure 2: EA dynamics for (a) tilted ferromagnetic states and (b) tilted antiferromagnetic states with varying $\gamma$.
 The blue curves correspond to $\theta=0.2\pi$, and the red curves represent $\theta=0.5\pi$.
-Panels (c) and (d) depict schematic 2-dimensional phase diagrams illustrating the dependence of early-time EA dynamics on $\theta$ and $1-\gamma$ for ferromagnetic and antiferromagnetic states, respectively.
+Panels (c) and (d) depict the dependence of early-time EA dynamics on $\theta$ and $1-\gamma$ for ferromagnetic and antiferromagnetic states, respectively.
 When the parameter is in the red region, EA can exceed the initial value, while in the blue region, EA firstly decreases and never grows higher than the initial value.
@@
 All calculations are based on $H_{1}$.
+Setup.— To study dynamics in these systems, we consider three initial states: the ferromagnetic state $|000...0\rangle$, the antiferromagnetic state $|0101..1\rangle$ and the domain-wall state $|000..111\rangle$, where the domain wall is positioned at the center of the chain.
+To incorporate the effect of symmetry breaking in the initial state, we introduce tilted ferromagnetic states [24, 25], defined as
+$|\psi_{i}(\theta)\rangle=e^{-i\frac{\theta}{2}\sum_{j}\sigma_{j}^{y}}|000...0\rangle$ (2)
+where $\sigma_{j}^{y}$ is the Pauli-$y$ matrix on the $j$-th qubit, and $\theta$ is a tuning parameter controlling the strength of symmetry breaking in the initial state.
+When $\theta=0$, Eq.
+(2) is U(1)-symmetric with zero EA.
+As $\theta$ increases, the EA grows, reaching its maximum value at $\theta=\pi/2$.
+The tilted antiferromagnetic and tilted domain wall states are constructed in a similar manner.
+The random circuit architecture consists of two-qubit U(1)-symmetric gates and Haar-random gates arranged in a brick-wall pattern.
+The U(1)-symmetric gates have a block-diagonal matrix structure, with each block independently sampled from the Haar measure [57, 58, 59].
+The effect of symmetry breaking depends on the density (doped probability) of random Haar gates without U(1) symmetry, denoted as $P_{\text{Haar}}$.
+The time unit in the circuit is defined by the application of two consecutive layers of gates.
+$\mathbb{E}[\Delta S_{a}]$ is computed by averaging over $5000$ circuit configurations.
+We also investigate Hamiltonian dynamics where the state $|\psi_{i}(\theta)\rangle$ undergoes unitary evolution by $e^{-iHt}|\psi_{i}(\theta)\rangle$, and the Hamiltonian is
+$H=$ $-\frac{1}{4}\sum_{j=1}^{L}\Big[\sigma_{j}^{x}\sigma_{j+1}^{x}+\gamma\sigma_{j}^{y}\sigma_{j+1}^{y}+\Delta_{1}\sigma_{j}^{z}\sigma_{j+1}^{z}\Big]$
+$-\Delta_{2}\sum_{j=1}^{L}\Big[\sigma_{j}^{x}\sigma_{j+2}^{x}+\sigma_{j}^{y}\sigma_{j+2}^{y}+\sigma_{j}^{z}\sigma_{j+2}^{z}\Big].$
+Here, $L$ denotes the total system size, while $\Delta_{1}$ and $\Delta_{2}$ represent the coefficients for nearest-neighbor and next-nearest-neighbor interactions, respectively.
+$\Delta_{2}$ introduces non-integrability, and $\gamma$ controls the strength of symmetry breaking.
+Periodic boundary conditions are imposed in both contexts.
+U(1)-Symmetric Initial States with U(1) Non-Symmetric Hamiltonian.— All numerical simulations are performed using the TensorCircuit-NG package [60].
+Here, we investigate the dynamics of symmetry breaking under an integrable Hamiltonian $H_{1}$ with $\Delta_{1}=0.4$ and $\Delta_{2}=0$, and a non-integrable Hamiltonian $H_{2}$ with $\Delta_{1}=0.4$ and $\Delta_{2}=0.05$, with system size $L=12$ sites.
+As revealed in Fig.
+1(a) and (b), EA for various different Hamiltonian symmetry-breaking $\gamma$ exhibit peaks at early times that are much larger than steady values.
+Furthermore, the peak value of the EA, $(\Delta S_{L/3})_{max}$, is found to be correlated with the strength of symmetry breaking, $1-\gamma$, for different symmetric initial states as shown in Fig.
+1(c) where EA of the ground state of $H_{1}$ follows the same trend.
+Notably, the peak heights nearly coincide between the ferromagnetic and domain wall states, as the early-time peak primarily depends on the local configurations of the initial state.
+Moreover, the finite-size scaling analysis demonstrates the existence of this peak in the thermodynamic limit (see SM).
+Ferromagnetic Domain Wall Antiferromagnetic
+EA(earlytime)EA\ (early\ time) crossing for small $1-\gamma$ crossing for small $1-\gamma$ crossing for small $1-\gamma$
+CV(earlytime)CV\ (early\ time) crossing for $\gamma\neq 1$ no crossing no crossing
+EA(latetime)EA\ (late\ time)
+CV(latetime)CV\ (late\ time)
+Table 1: The early- and late-time behavior of EA and CV under the evolution of $H_{1}$ or $H_{2}$ ($0.5\leq\gamma\leq 1$).
+Crossing in EA (CV) means when the time-evolution curves of EA (CV) for states with larger $\theta$ intersect with those for smaller $\theta$ at early-times.
+The right-up (right-down) arrow indicates that the late-time value is increasing (decreasing) with increasing tilted angle $\theta$.
 By analyzing Fig.
-3, we identify that the late-time EA, denoted as $\Delta S_{L/3}^{\infty}$, oscillates and does not approach zero.
+1, we identify that the late-time EA $\Delta S_{L/3}^{\infty}$ oscillates and does not approach zero.
 This is because the reduced density matrix of subsystem $a$ evolves towards a canonical ensemble $e^{-\beta\hat{H_{a}}}$, where $\hat{H_{a}}$ has the same form as $\hat{H}$ in Eq.
 (Symmetry Breaking Dynamics in Quantum Many-Body Systems), but acts solely on subsystem $a$.
-Since $\hat{H_{a}}$ includes symmetry breaking terms, $[\rho_{a},\hat{Q_{a}}]\neq 0$, leading to a non-vanishing EA at long times.
+Since $\hat{H_{a}}$ includes symmetry breaking terms, $\rho_{a},\hat{Q_{a}}]\neq 0$, leading to a non-vanishing EA at long times.
 In Fig.
-3 (d), we calculate the ratio of $\Delta S_{L/3}^{\infty}$ to $(\Delta S_{L/3})_{max}$ with varying $\gamma$ for different initial states.
+1(d), we calculate the ratio of $\Delta S_{L/3}^{\infty}$ to $(\Delta S_{L/3})_{max}$ with varying $\gamma$.
 The late-time EA, $\Delta S_{L/3}^{\infty}$, is obtained by averaging $\Delta S_{L/3}$ over $2000$ time points between $t_{1}=2000$ and $t_{2}=40000$.
-The results further confirm the overshooting behavior as the late-time saturating EA value is much lower than the peak value at the early time.
-On the contrary, the CV dynamics in this setting shows no evident overshooting pattern but instead directly grows to the saturating values (see SM).
-U(1)-Asymmetric Initial States with U(1) Non-Symmetric Hamiltonian.— Next we investigate the behavior of EA with U(1)-asymmetric initial states under $H_{1}$.
-Now, the time evolution of EA depends on both symmetry-breaking parameters, $\theta$ and $\gamma$.
-$\theta$ describes the symmetry breaking in the initial state and determines the initial value of EA.
-$\gamma$, on the other hand, serves as a measure of symmetry breaking in the Hamiltonian, which influences the evolution of the state.
-The interplay between these two parameters results in distinct behaviors in the time dependence of EA.
-This is illustrated in the schematic 2-dimensional phase diagrams with varying $\theta$ and $\gamma$ in Fig.
-4 (c) and (d).
+The results further confirm the overshooting behavior as the late-time saturating EA value is much lower than the early-time peak value.
+This phenomenon stems from the competition between symmetry breaking and subsystem decoherence.
+Initially, the non-symmetric Hamiltonian dynamically generates asymmetry, causing EA to grow as symmetry breaking predominates over decoherence.
+Later, symmetry breaking saturates while decoherence becomes the sole governing factor for EA evolution, driving EA toward a steady-state value.
+The early-time peak emerges from the intricate interplay between these competing effects.
+On the contrary, the CV dynamics in this setting shows no evident overshooting pattern but instead directly grows to the saturating values.
+Figure 3: (a) Schematic illustration of a non-symmetric random circuit with 6 qubits.
+Gates are arranged in the even-odd brick-wall pattern.
+The blue and red rectangles represent U(1)-symmetric and random Haar gates, respectively.
+The basis for the U(1)-symmetric gate is listed in the following order: $|00\rangle$, $\rangle, |10|10\rangle$ and $|11\rangle$.
+(b) The circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]$, as a function of time with the antiferromagnetic initial state at different values of $P_{\text{Haar}}$.
+(c) The peak value, $\mathbb{E}[\Delta S_{L/4}]_{max}$, as a function of $P_{\text{Haar}}$.
+All three curves follow a power law $y=ax^{b}$.
+F: Ferromagnetic state ($a=1.4$, $b=0.4$); DW: Domain Wall state ($a=2.7$, $b=0.8$); AF: Antiferromagnetic state ($a=1.9$, $b=0.9$).
+U(1)-Asymmetric Initial States with U(1) Non-Symmetric Hamiltonian.— Next we investigate dynamics with U(1)-asymmetric initial states under $H_{1}$ where EA dynamics depends on both symmetry-breaking parameters, $\theta$ and $\gamma$.
+$\theta$ describes the symmetry breaking in the initial state while $\gamma$ describes the symmetry breaking in the Hamiltonian.
+The interplay between these two parameters results in distinct behaviors in the EA dynamics.
+This is illustrated in the schematic figures with varying $\theta$ and $\gamma$ in Fig.
+2(c) and (d).
 The colors highlight the tendency in EA at early times.
 Blue regime indicates that $\Delta S_{L/4}(t)$ never exceeds its initial value for early times, while red regime corresponds to the situations where EA can grow larger than its initial value at early times.
-It is clearly reflected in Figs.
-4 (a), the initial growth of EA at $\theta=0.2\pi$ and $\gamma=0.8$, $0.6$ aligns with the red region shown in the phase diagram.
-Two key observations are made from the phase diagram: (1) For a fixed $\gamma$, EA grows with weaker asymmetric effects (small $\theta$) in the initial states or for a fixed $\theta$, EA increases with stronger symmetry breaking effects (large $1-\gamma$) in the Hamiltonian.
-(2) The threshold for $\gamma$ at which EA begins to rise varies with different initial states.
-As illustrated in Fig.
-4 (d), the range in which EA exceeds the initial value is very limited with an antiferromagnetic initial state.
-In other words, the early-time behavior of EA serves as a witness to compare the symmetry-breaking strength hosting in the quantum state and the Hamiltonian.
-When the Hamiltonian is more asymmetric than the initial state, EA will first increase, often resulting in the overshooting behavior in the early time, as indicated by the red regime in the phase diagram.
+It is clearly reflected in Fig.
+2(a), the initial growth of EA at $\theta=0.2\pi$ and $\gamma=0.8$, $0.6$ aligns with the red region.
+For a fixed $\gamma$, EA grows with weaker asymmetric effects (small $\theta$) in the initial states or for a fixed $\theta$, EA increases with stronger symmetry breaking effects (large $1-\gamma$) in the Hamiltonian.
+Consequently, the early-time behavior of EA serves as a witness to compare the symmetry-breaking strength hosting in the quantum state and the Hamiltonian.
 Another key feature of the early-time dynamics is the emergence of QME, as shown in Fig.
-4 (a) for the symmetric case $\gamma=1$.
+2(a) for the symmetric case $\gamma=1$.
 The origin of this QME lies in the relative small $ZZ$ term and gapless nature in the Hamiltonian [42].
-Notably, QME persists for ferromagnetic (antiferromagnetic) states when $0.8\leq\gamma\leq 1$ ($0.4\leq\gamma\leq 1$).
-The robustness of QME against weak symmetry-breaking is a general feature in quantum many-body systems when the Hamiltonian is non-symmetric.
+QME persists for ferromagnetic (antiferromagnetic) states when $0.8\leq\gamma\leq 1$ ($0.4\leq\gamma\leq 1$).
 We also report relevant results for non-integrable Hamiltonian $H_{2}$ in the SM, and the results remain qualitatively consistent with cases of $H_{1}$, demonstrating the universal applicability of conclusions in this Letter for Hamiltonian evolutions.
-Our simulation on another symmetry-breaking measure, charge variance, reveals that QME can also emerge for CV with initial tilted ferromagnetic states, but only in cases of non-symmetric evolution.
+Our simulation on the other symmetry-breaking measure, charge variance, reveals that QME can also emerge for CV with initial tilted ferromagnetic states, but only in cases of non-symmetric evolution.
 The reversed monotonicity of CV with respect to $\theta$ can persist even at late times.
@@
 The distinction shows the richness in characterizing symmetry breaking strength and patterns.
-Ferromagnetic Domain Wall Antiferromagnetic
-$EA\ (early\ time)$ crossing for small $1-\gamma$ crossing for small $1-\gamma$ crossing for small $1-\gamma$
-$CV\ (early\ time)$ crossing for $\gamma\neq 1$ no crossing no crossing
-$EA\ (late\ time)$
-$CV\ (late\ time)$
-Table 1: The early- and late-time behavior of EA and CV under the evolution of $H_{1}$ or $H_{2}$ ($0.5\leq\gamma\leq 1$).
-The crossing observed at early times signals the occurrence of the QME.
-The right-up (right-down) arrow indicates that the late-time value is increasing (decreasing) with increasing tilted angle $\theta$.
-Conclusions and discussions.— In this Letter, we present a comprehensive study of subsystem symmetry breaking within two frameworks: a non-symmetric random circuit and a non-symmetric Hamiltonian evolution.
+U(1)-Symmetric (Asymmetric) States with U(1) Non-Symmetric Random Circuit.— A schematic diagram of the circuit architecture is shown in Fig.
+3(a).
+The circuit under consideration consists of 16 qubits.
+We evaluate the EA at different $P_{\text{Haar}}$, using an antiferromagnetic initial state.
+We observe that EAs approach zero at late times, as illustrated in Fig.
+3(b).
+This behavior can be understood in the context of quantum thermalization and information scrambling [61, 62, 63, 64], where the reduced density matrix of the subsystem is a fully mixed state for the random circuit cases, as long as the subsystem size does not exceed half of the total system.
+Additionally, for all probabilities chosen in Fig.
+3(b), EAs reach their maximum after only a few layers of unitaries.
+The rate of symmetry restoration also depends on the initial state.
+In the SM, we find that symmetry restoration occurs more quickly for antiferromagnetic or domain wall states than for ferromagnetic states, due to the larger Hilbert space sector of the initial states in the former cases.
+In Fig.
+3(c), we reveal that the peak of the circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]_{max}$ follows a power-law with respect to $P_{\text{Haar}}$ for small $P_{\text{Haar}}$.
+Next, we examine the dynamics from U(1)-asymmetric initial states, i.e.
+a tilted ferromagnetic state.
+We compute the EA for both U(1) symmetry with $\hat{Q_{a}}=\sum_{i\in a}\sigma_{i}^{z}$ and $Z_{2}$ symmetry with $\hat{Q_{a}}=\prod_{i\in a}\sigma_{i}^{z}$.
+As depicted in Fig.
+4(a), for $P_{\text{Haar}}=0$, we clearly notice the emergence of QME in U(1) case.
+Surprisingly, we also find that the QME appears in the $Z_{2}$ probe, which does not contradict previous study [25] suggesting the absence of QME in $Z_{2}$-symmetric circuits.
+Even though U(1)-symmetric gates are also $Z_{2}$ symmetric, there is no off-diagonal coupling between $|00\rangle$ and $|11\rangle$, leading to different thermalization rates between two $Z_{2}$ charge sectors ($Q_{a}=\pm 1$), and thus resulting in QME.
+As we replace a portion of U(1)-symmetric gates with random Haar gates, QME remains evident with a finite number of random Haar gates.
+However, when the circuit consists entirely of random Haar gates, all charge sectors thermalize at the same rate after circuit averaging, and QME disappears.
+In this case, the overshooting mechanism becomes apparent: EA curves for all initial states (parametrized by $\theta$) converge to a common finite value after a single gate layer, then decay toward zero at late times.
+Overshooting occurs when this transient value exceeds the initial EA of certain states - particularly symmetric initial states.
+Conclusions and discussions.— In this Letter, we present a comprehensive study of subsystem symmetry breaking within two frameworks: a non-symmetric Hamiltonian evolution and a non-symmetric random circuit.
 Our simulation reveals that U(1) symmetry is always restored in the non-symmetric random circuit case, regardless of the initial states or the density of symmetry-breaking random Haar gates $P_{\text{Haar}}$.
 On the contrary, subsystem U(1) symmetry remains broken in the case of a U(1) non-symmetric Hamiltonian.
+Figure 4: The circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]$, as a function of time for different values of $P_{\text{Haar}}$.
+Blue: U(1) EA.
+Green: $Z_{2}$ EA.
+Panels (a)-(d) correspond to different values of $P_{\text{Haar}}$ (a) $P_{\text{Haar}}=0$, (b) $P_{\text{Haar}}=0.3$, (c) $P_{\text{Haar}}=0.7$, and (d) $P_{\text{Haar}}=1$, respectively.
 In addition to the late-time results, the early-time dynamics of EA shows a universal and surprising feature of overshooting.
@@
 Data availability. Numerical data for this manuscript are publicly accessible in Ref. [91].
-Supplemental Material for “Symmetry Breaking Dynamics in Quantum Many-Body Systems”
-February 8, 2025
-I More numerical results for U(1) non-symmetric quantum circuits
-I.1 Dynamics of EA with other U(1)-symmetric initial states
-In this section, we present numerical results of the dynamics of entanglement asymmetry with different initial states, using the same setup as described in the main text.
-Similar to the behavior observed in Fig.
-1 (b), EA for the other initial states in Fig.
-S1 also exhibits an initial peak, which then gradually decays to zero, indicating the restoration of U(1) symmetry.
-It is most apparent from the case of ferromagnetic states that a smaller density of random Haar gates results in a longer time of symmetry restoration in the subsystem.
-In addition, the thermalization process occurs more rapidly in domain wall and antiferromagnetic states than ferromagnetic states.
-This is because, in the limit of symmetry-preserving circuits, the Hilbert space accessible to the antiferromagnetic states is much larger than the ferromagnetic cases.
-Figure S1: The dynamics of circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]$ for different initial states with $L=16$.
-The varying color intensity represents different densities of random Haar gates.
-Left: Ferromagnetic state.
-Right: Domain Wall state.
-I.2 Dynamics of EA with U(1)-asymmetric initial states at different $P_{\text{Haar}}$
-In addition to the $P_{\text{Haar}}$ values used in the main text, we show the behavior of entanglement asymmetry with tilted ferromagnetic state for other $P_{\text{Haar}}$ values in Fig.
-S2, and find that QME is indeed present except when the symmetry breaking reaches the maximum $P_{\text{Haar}}=1$.
-This phenomenon also holds true for tilted domain wall state, as demonstrated in Fig.
-S3.
-On the contrary, as shown in Fig.
-S4, QME is absent for tilted antiferromagnetic states across all values of $P_{\text{Haar}}$.
-Therefore, the emergence of QME is specific to certain initial states and is robust against the effects of random Haar gates on the circuit evolution.
-Figure S2: The dynamics of circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]$, for tilted ferromagnetic initial states is examined for various values of $P_{\text{Haar}}$ with $L=16$.
-Blue: U(1) EA.
-Green: $Z_{2}$ EA.
-Panels (a)-(d) correspond to the following values of $P_{\text{Haar}}$ (a) $P_{\text{Haar}}=0.1$, (b) $P_{\text{Haar}}=0.2$, (c) $P_{\text{Haar}}=0.5$, and (d) $P_{\text{Haar}}=0.9$.
-Figure S3: The dynamics of circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]$, for tilted domain wall initial states is examined for various values of $P_{\text{Haar}}$ with $L=16$.
-Blue: U(1) EA.
-Green: $Z_{2}$ EA.
-Panels (a)-(d) correspond to the following values of $P_{\text{Haar}}$ (a) $P_{\text{Haar}}=0$, (b) $P_{\text{Haar}}=0.3$, (c) $P_{\text{Haar}}=0.7$, and (d) $P_{\text{Haar}}=1$.
-Figure S4: The dynamics of circuit-averaged EA, $\mathbb{E}[\Delta S_{L/4}]$, for tilted antiferromagnetic initial states is examined for various values of $P_{\text{Haar}}$ with $L=16$.
-Blue: U(1) EA.
-Green: $Z_{2}$ EA.
-Panels (a)-(d) correspond to the following values of $P_{\text{Haar}}$ (a) $P_{\text{Haar}}=0$, (b) $P_{\text{Haar}}=0.3$, (c) $P_{\text{Haar}}=0.7$, and (d) $P_{\text{Haar}}=1$.
-I.3 Dynamics of CV with U(1)-symmetric(asymmetric) initial states at different $P_{\text{Haar}}$
-As shown in Fig.
-S5, the circuit-averaged charge variance $\mathbb{E}[\sigma^{2}_{Q}]$ for all U(1)-symmetric initial states gradually increases until it reaches a plateau.
-This behavior contrasts with the overshooting dynamics observed in the entanglement asymmetry.
-Besides, the early-time dynamics of charge variance from different asymmetric initial states show no evidence of QME, as confirmed in Fig.
-S6.
-Two observations can be drawn from Figs.
-S5 and S6: (1) The time required to reach the plateau decreases as the density of random Haar gates increases.
-Furthermore, the charge variance saturates more quickly for the antiferromagnetic state compared to the ferromagnetic state.
-(2) The charge variance saturates to the same value for all U(1)-symmetric(asymmetric) initial states.
-This arises because the late-time density matrix for a small subsystem approaches an identity matrix.
-Here, $\mathbb{E}[\sigma^{2}_{Q}]$ is computed by averaging the charge variance, $\sigma^{2}_{Q}$, over $5000$ circuit configurations.
-Figure S5: The dynamics of circuit-averaged CV, $\mathbb{E}[\sigma^{2}_{Q}]$, for different U(1)-symmetric initial states are examined for various values of $P_{\text{Haar}}$ with $L=16$.
-From left to right: the initial states are ferromagnetic, domain wall and antiferromagnetic states, respectively.
-Figure S6: Dynamics of circuit averaged CV, $\mathbb{E}[\sigma^{2}_{Q}]$, for different U(1)-asymmetric initial states in a system with $L=16$.
-The top, middle, and bottom rows correspond to tilted ferromagnetic, tilted domain wall, and tilted antiferromagnetic initial states, respectively.
-The left and columns show results for $P_{\text{Haar}}=0.05$ and $P_{\text{Haar}}=0.1$, respectively.
-I.4 The early and late time behavior of EA and CV in U(1) non-symmetric quantum circuits
-In Table.
-2, we provide a summary of the behavior of entanglement asymmetry and charge variance at early and late times for U(1)-asymmetric states in the circuit model.
-We observe that the quantum Mpemba effect occurs only in the early-time dynamics of entanglement asymmetry for tilted ferromagnetic and tilted domain wall states.
-Moreover, the entanglement asymmetry and charge variance approach distinct values, which are universal across all initial states, at late times.
-Ferromagnetic Domain Wall Antiferromagnetic
-$EA\ (early\ time)$ crossing crossing no crossing
-$CV\ (early\ time)$ no crossing no crossing no crossing
-$EA\ (late\ time)$ 0 0 0
-$CV\ (late\ time)$ C C C
-Table 2: The early- and late-time behavior of entanglement asymmetry (EA) and charge variance (CV) under the evolution of random circuit with $10\leq P_{\text{Haar}}<$ 1.
-The crossing in the behavior indicates the occurrence of the quantum Mpemba effect.
-The constant C can be evaluated as $\mathrm{tr}(\rho Q^{2})-\mathrm{tr}(\rho Q)^{2}$, where $\rho=\frac{I}{2^{L}}$ is the late-time density matrix for the subsystem.
-II More numerical results of entanglement asymmetry for U(1) non-symmetric Hamiltonians
-II.1 Dynamics of entanglement asymmetry for various U(1)-symmetric initial states
-In Fig.
-S7, we show the dynamics of EA across all initial states (ferromagnetic, domain wall, and antiferromagnetic) under the evolution of $H_{2}$.
-The entanglement asymmetry shows evident overshooting at early times, and the U(1) symmetry remains broken in the subsystem at long times in all cases, consistent with the results described in the main text.
-As shown in Fig.
-S8, the peak of entanglement asymmetry also correlates with the strength of symmetry breaking, $1-\gamma$, for various initial states in the case of a non-integrable Hamiltonian.
-EA for ground state of $H_{2}$ is also shown for comparison.
-More importantly, the peak value of entanglement asymmetry is significantly higher than its late-time value.
-Figure S7: EA as a function of time for various initial states and values of $\gamma$, with $L=12$.
-Panels (a) are based on the $H_{1}$, while panels (b), (c), and (d) correspond to $H_{2}$.
-Insets zoom in on the peak of EA for different values of $\gamma$ at early times, listed from bottom to top as: $\gamma=0.8,0.7,0.6,0.5,0.4,0.3,0.2,0.1$.
-(a),(c): Domain Wall state.
-(b): Ferromagnetic state.
-(d): Antiferromagnetic state.
-Figure S8: Panels (a) and (b) show the peak value of EA, $(\Delta S_{L/3})_{max}$, and the ratio of the late-time EA, $\Delta S_{L/3}^{\infty}$, to $(\Delta S_{L/3})_{max}$ as a function of $1-\gamma$ for different initial states under $H_{2}$.
-F: Ferromagnetic state; DW: Domain Wall state; AF: Antiferromagnetic state; GS: Ground State.
-II.2 Dynamics of entanglement entropy
-To better explain the behavior of entanglement asymmetry, we examine the time evolution of entanglement entropy and explore its connection to entanglement asymmetry in this section.
-We present both $S_{L/3}$ and $S_{L/3,Q}$ as functions of time, with their difference yielding EA, for various initial states and Hamiltonians.
-As shown in Fig.
-S9, we observe that the entanglement entropy grows linearly with time before eventually reaching a plateau.
-The saturation time of the entanglement entropy is always greater than the time at which the entanglement asymmetry reaches its peak.
-In most cases, the peak of entanglement asymmetry occurs after $S_{L/3,Q}$ reaches its first peak.
-This observation is reflected in Fig.
-S9.
-Figure S9: Dynamics of each individual term in the expression for the entanglement asymmetry $\Delta S_{L/3}$ with $L=12$ is shown.
-Solid lines represent $S_{L/3,Q}$, while dash lines correspond to the entanglement entropy $S_{L/3}$.
-From left to right: initial states are Ferromagnetic, Domain wall and Antiferromagnetic states, respectively.
-Panels (a), (b), and (c) are obtained using $H_{1}$, while panels (d), (e), and (f) are based on $H_{2}$.
-II.3 Quantum Mpemba effect in U(1) non-symmetric Hamiltonian
-When $\gamma=1$, the emergence of QME is observed in all cases, as shown in Figs.
-S10 and S11.
-Analogous to the discussion in the main text, we find the absence of QME in the non-integrable Hamiltonian across all initial states when $\gamma$ is tuned to above some thresholds.
-Figure S10: EA dynamics at different values of $\gamma$ with $L=12$ are examined for tilted domain wall states.
-The calculation is based on $H_{1}$.
-Blue corresponds to $\theta=0.2\pi$, and red corresponds to $\theta=0.5\pi$.
-The increasing intensity of color reflects a stronger effect of symmetry breaking.
-Figure S11: The dynamics of entanglement asymmetry are examined for different values of $\gamma$ with $L=12$ across three initial states: (a) tilted ferromagnetic states, (b) tilted domain wall states, and (c) tilted antiferromagnetic states.
-Panels (a)–(c) are based on the evolution under $H_{2}$.
-The blue curves correspond to $\theta=0.2\pi$, while the red curves correspond to $\theta=0.5\pi$.
-The increasing intensity of the color reflects the stronger effect of symmetry breaking.
-II.4 Phase diagrams in U(1) non-symmetric Hamiltonians
-As shown in Figs.
-S12 and S13, the phase diagrams for ferromagnetic and antiferromagnetic states obtained through $H_{2}$ are almost identical to those from $H_{1}$.
-A slight difference is observed in the case of the domain wall state.
-Since the phase boundaries are different for different initial states, a natural question arises – is the phase boundary only different in terms of $\theta$ or also different when translating $\theta$ to initial EA.
-In Fig.
-S14, we replace $\theta$ with the initial value of the entanglement asymmetry, $\Delta S_{L/4}^{ini}$, on the $y$-axis.
-We find that the phase boundary is still different in terms of initial EA for different types of initial states.
-Figure S12: A schematic phase diagram illustrating the dependence of EA dynamical patterns on the parameters $\theta$ and $1-\gamma$ for a domain wall state under the evolution of $H_{1}$.
-Those black dots are obtained through numerical simulation.
-Figure S13: Panels (a)-(c) show schematic phase diagrams depicting the dependence of EA dynamical patterns on $\theta$ and $1-\gamma$ for different initial states.
-All panels are based on $H_{2}$.
-(a) Ferromagnetic state, (b) Domain Wall state, (c) Antiferromagnetic state.
-Those black dots are obtained through numerical simulation.
-Figure S14: Panels (a)-(f) show schematic phase diagrams depicting the dependence of EA dynamical patterns on the initial value of EA, $\Delta S_{L/4}^{ini}$ and $1-\gamma$ for different initial states and Hamiltonians.
-From left to right: initial states are Ferromagnetic, Domain wall and Antiferromagnetic states, respectively.
-Panels (a), (b), and (c) are obtained using $H_{1}$, while panels (d), (e), and (f) are based on $H_{2}$.
-II.5 Late-time entanglement asymmetry
-The late-time entanglement asymmetry is obtained by averaging $\Delta S_{L/4}$ over $2000$ time points between $t_{1}=2000$ and $t_{2}=40000$, as described in the main text.
-In S15, we compute the late-time EA for different initial states under the evolution of $H_{1}$ and $H_{2}$.
-We find two general trends: (1) The late-time value of EA increases with the strength of symmetry breaking $1-\gamma$.
-(2) For states with higher initial asymmetry, the entanglement asymmetry at late times is greater and increases more rapidly compared to states with lower initial asymmetry.
-Figure S15: The late-time entanglement asymmetry, $\Delta S_{L/4}^{\infty}$, as a function of $1-\gamma$ with $L=12$.
-From left to right: the initial states are tilted ferromagnetic, tilted domain wall and tilted antiferromagnetic states, respectively.
-Panels (a), (b), and (c) correspond to results obtained using $H_{1}$, while panels (d), (e), and (f) are based on $H_{2}$.
-III More numerical results of charge variance for U(1) non-symmetric Hamiltonians
-Figure S16: The time evolution of the expectation value of the total spin in the $z$-direction, $\langle Q_{tot}\rangle$, is investigated for various values of $\gamma$ with a ferromagnetic initial state, where $L=12$.
-Here, $\hat{Q}_{tot}=\sum_{i=1}^{L}\sigma_{i}^{z}$.
-Left: $H_{1}$.
-Right: $H_{2}$.
-Figure S17: The variance of $Q_{tot}$, $\sigma_{Q}^{2}$, as a function of time for different values of $\gamma$ with $L=12$.
-Top row: $H_{1}$.
-Bottom row: $H_{2}$.
-From left to right: Ferromagnetic, Domain Wall, and Antiferromagnetic states.
-III.1 Expectation values of $\hat{Q}$ and $\sigma_{Q}^{2}$
-The expectation value of the total spin in the $z$-direction as a function of time is shown in Fig.
-S16 for ferromagnetic states.
-At the start, $Q_{tot}=L=12$ since all the spins are aligned upward.
-Over time, the state progressively loses memory of its initial configuration, and $Q_{tot}$ settles into oscillations around 00.
-For domain wall and antiferromagnetic states, $\langle\hat{Q}_{tot}(t)\rangle=0$ is a constant due to symmetry argument which are elaborated below.
-Suppose we have a transformation $T$:
-$T:$ $\quad\sigma^{z}\rightarrow-\sigma^{z},\quad\sigma^{y}\rightarrow-\sigma^{y}$ (S1)
-It is straightforward to verify that the Hamiltonians remain invariant under this transformation $T$. Thus, we have
-$\langle\hat{Q}_{\text{tot}}(t)\rangle$ $=\langle\psi(t)|\hat{Q}_{\text{tot}}|\psi(t)\rangle$
-$=\langle\psi(0)|e^{iHt}\hat{Q}_{\text{tot}}e^{-iHt}|\psi(0)\rangle$
-$=\langle\psi(0)|T^{\dagger}Te^{iHt}T^{\dagger}T\hat{Q}_{\text{tot}}T^{\dagger}Te^{-iHt}T^{\dagger}T|\psi(0)\rangle$
-$=\langle\psi(0)|T^{\dagger}e^{iHt}(-\hat{Q}_{\text{tot}})e^{-iHt}T|\psi(0)\rangle$ (S2)
-$=\langle\psi_{p}(0)|e^{iHt}(-\hat{Q}_{\text{tot}})e^{-iHt}|\psi_{p}(0)\rangle$
-$=-\langle\psi_{p}(t)|\hat{Q}_{\text{tot}}|\psi_{p}(t)\rangle$
-$=-\langle\psi(t)|\hat{Q}_{\text{tot}}|\psi(t)\rangle$
-$=0$
-Here $|\psi_{p}(t)\rangle$ is related to $|\psi(t)\rangle$ by a permutation of site indices.
-Throughout the derivation, we use the fact that $T\hat{Q}_{\text{tot}}T^{\dagger}=-\hat{Q}_{\text{tot}}$.
-Therefore, $\langle\hat{Q}_{tot}(t)\rangle$ is strictly zero for domain wall and antiferromagnetic initial states.
-Fig.
-S17 shows the time dependence of the variance of total charge operator for different initial states under $H_{1}$ and $H_{2}$.
-Notably, the variance of the charge typically saturates later than the entanglement asymmetry.
-The charge variance also characterizes the U(1) symmetry breaking in the evolved state from some aspects, similar as EA explored in the main text.
-However, the symmetry breaking dynamical behaviors show distinct patterns in the two metrics.
-EA shows an evident overshooting while the charge variance directly saturates.
-The differences may provide further insight into the multifaceted physics of symmetry breaking.
-III.2 Dynamics of charge variance for different U(1)-asymmetric initial states
-Figures S18 and S19 show that, for charge variance evolved from the tilted domain wall and tilted antiferromagnetic states, the charge variance is consistently larger for more asymmetric states (large $\theta$) than for less asymmetric ones (small $\theta$).
-However, this trend does not apply to the initial tilted ferromagnetic state, where the monotonic relationship of charge variance with respect to $\theta$ is reversed at early times.
-This results in an early-time crossing, which becomes more pronounced as $1-\gamma$ increases.
-These behaviors are observed in both integrable and non-integrable Hamiltonians, highlighting their universality across different dynamical regimes.
-Figure S18: Temporal evolution of the charge variance, $\sigma_{Q}^{2}$, under the integrable Hamiltonian $H_{1}$ for a system size $L=12$.
-The top row corresponds to tilted ferromagnetic states, the middle row to tilted domain wall states, and the bottom row to tilted antiferromagnetic states.
-From left to right, the columns show results for $\gamma=0.9$, $0.8$, and $0.7$.
-Figure S19: Time evolution of the charge variance, $\sigma_{Q}^{2}$, under the non-integrable Hamiltonian $H_{2}$ for a system size $L=12$.
-The top row corresponds to tilted ferromagnetic states, the middle row to tilted domain wall states, and the bottom row to tilted antiferromagnetic states.
-From left to right, the columns show results for $\gamma=0.9$, $0.8$, and $0.7$.
-III.3 Dynamics of charge sector probability distributions
-Figure S20 displays the probability distribution $P_{Q}$ across various charge sectors $Q$ for the evolved state $|\psi(t)\rangle$ at different times for initial tilted ferromagnetic states.
-Here, $P_{Q}$ at time $t$ is defined as $\sum_{q=Q}|\langle\psi(t)|\psi_{q}\rangle|^{2}$, where $|\psi_{q}\rangle$ represents the basis wave function corresponding to charge $q$.
-For more asymmetric initial states (larger $\theta$), the charge distribution shrinks, leading to a decrease in charge variance.
-In contrast, for less asymmetric states, the distribution spreads across a broader range of charge sectors, increasing the charge variance.
-This behavior accounts for the crossing phenomenon observed in Figs.
-S18 and S19.
-Figure S20: Time evolution of the probability distribution, $P_{Q}$, for each charge sector $Q$ under the Hamiltonian $H_{1}$ with $\gamma=0.6$.
-The charge sectors range from $Q=-12$ to $Q=12$ in increments of $2$.
-The rows, from top to bottom, correspond to time points $t=0$, $0.75$, $1.5$, $3$, and $4$.
-The columns, from left to right, represent different tilted ferromagnetic states with $\theta=0.2\pi$, $0.3\pi$, $0.4\pi$, and $0.5\pi$.
-III.4 Phase diagrams in U(1) non-symmetric Hamiltonian
-Figure S21 presents schematic 2-dimensional phase diagrams that capture the dependence of early-time CV dynamics on the parameters $\theta$ and $1-\gamma$ for the ferromagnetic state evolving under $H_{1}$ and $H_{2}$.
-The diagram derived from $H_{2}$ closely resembles those obtained from $H_{1}$.
-The red region indicates where CV initially grows, While the blue region corresponds to an initial decrease in CV.
-At $\gamma=1$ (left boundary), the charge variance remains constant and equal to its initial value.
-Figure S21: Panel (a) and (b) present schematic phase diagrams depicting the dependence of charge variance on $\theta$ and $1-\gamma$ with ferromagnetic state evolving under $H_{1}$ (left) and $H_{2}$ (right).
-III.5 Late-time charge variance
-The late-time charge variance, $\sigma_{Q}^{2}(t\rightarrow\infty)$, is computed in a manner similar to the late-time entanglement asymmetry.
-We examine the late-time charge variance for different initial states under the evolution of both $H_{1}$ and $H_{2}$, as illustrated in Fig.
-S22.
-It is observed that the late-time charge variance increases with the tilted angle $\theta$ for domain wall and antiferromagnetic states at a fixed $\gamma$.
-In contrast, for ferromagnetic states, the late-time charge variance decreases as $\theta$ increases, which is highly non-trivial.
-This behavior, also observed at early times (as seen in the crossing in Figs.
-S18 and S19), continues into the late-time regime.
-Figure S22: The late-time charge variance, $\sigma_{Q\infty}^{2}$, as a function of $1-\gamma$ with $L=12$.
-From left to right: the initial states are tilted ferromagnetic, tilted domain wall and tilted antiferromagnetic states, respectively.
-Panels (a), (b), and (c) correspond to results obtained using $H_{1}$, while panels (d), (e), and (f) are based on $H_{2}$.
-IV Analytical results of charge variance at early times
-The early-time dynamics of charge variance, $\sigma_{Q}^{2}=\langle Q^{2}\rangle-\langle Q\rangle^{2}$, can be expanded as:
-$\sigma_{Q}^{2}(t)\approx\sigma_{Q}^{2}(0)+t\left.\frac{d\sigma_{Q}^{2}}{dt}\right|_{0}+\frac{t^{2}}{2}\left.\frac{d^{2}\sigma_{Q}^{2}}{dt^{2}}\right|_{0}+...$ (S3)
-where the ellipsis represents higher order terms that are neglected.
-Here, $\left.\frac{d\sigma_{Q}^{2}}{dt}\right|_{0}$ and $\left.\frac{d^{2}\sigma_{Q}^{2}}{dt^{2}}\right|_{0}$ denote the first and second derivatives of the charge variance evaluated at $t=0$, respectively.
-The initial state is chosen to be a tilted ferromagnetic state.
-Using the Heisenberg equation of motion, the first derivative of the charge variance at $t=0$ is given by:
-$\left.\frac{d\sigma_{Q}^{2}}{dt}\right|_{0}=i\langle[H,Q^{2}]\rangle_{0}-2i\langle Q\rangle_{0}\langle[H,Q]\rangle_{0}$ (S4)
-To evaluate $[H,Q]$, where the Hamiltonian is:
-$H=-\frac{1}{4}\sum_{i=1}^{L}(\sigma_{i}^{x}\sigma_{i+1}^{x}+\gamma\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta_{1}\sigma_{i}^{z}\sigma_{i+1}^{z})$ (S5)
-and the charge operator is $Q=\sum_{j=1}^{L}\sigma_{j}^{z}$, We compute:
-$[H,Q]$ $=$ $[-\frac{1}{4}\sum_{i=1}^{L}(\sigma_{i}^{x}\sigma_{i+1}^{x}+\gamma\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta_{1}\sigma_{i}^{z}\sigma_{i+1}^{z}),\sum_{j=1}^{L}\sigma_{j}^{z}]$ (S6)
-$=$ $-\frac{1}{4}\sum_{j}[\sigma_{j-1}^{x}\sigma_{j}^{x}+\sigma_{j}^{x}\sigma_{j+1}^{x}+\gamma\sigma_{j-1}^{y}\sigma_{j}^{y}+\gamma\sigma_{j}^{y}\sigma_{j+1}^{y},\sigma_{j}^{z}]$ (S7)
-Simplifying, we obtain:
-$[H,Q]=\frac{i}{2}(1-\gamma)\sum_{j}(\sigma_{j}^{x}\sigma_{j+1}^{y}+\sigma_{j}^{y}\sigma_{j+1}^{x})$ (S8)
-Since $\langle\sigma_{j}^{y}\rangle_{0}=0$, it follows that $\langle[H,Q]\rangle_{0}=0$.
-Similarly, one can show that $\langle[H,Q^{2}]\rangle_{0}=0$.
-Consequently, the linear term in the expansion Eq.
-S3 vanishes.
-The quadratic term in the expansion is determined by the second derivative of the charge variance:
-$\left.\frac{d^{2}\sigma_{Q}^{2}}{dt^{2}}\right|_{0}=-\langle[H,[H,Q^{2}]]\rangle_{0}+2\langle Q\rangle_{0}\langle[H,[H,Q]]\rangle_{0}$ (S9)
-To compute $[H,[H,Q]]$, we evaluate:
-$[H,[H,Q]]$ $=$ $[-\frac{1}{4}\sum_{i=1}^{L}(\sigma_{i}^{x}\sigma_{i+1}^{x}+\gamma\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta_{1}\sigma_{i}^{z}\sigma_{i+1}^{z}),\frac{i}{2}(1-\gamma)\sum_{j}(\sigma_{j}^{x}\sigma_{j+1}^{y}+\sigma_{j}^{y}\sigma_{j+1}^{x})]$ (S10)
-$=$ $\sum_{j}\frac{1}{2}(1-\gamma)((1-\gamma)\sigma_{j}^{z}+\sigma_{j-1}^{x}\sigma_{j}^{z}\sigma_{j+1}^{x}-\gamma\sigma_{j-1}^{y}\sigma_{j}^{z}\sigma_{j+1}^{y})$ (S11)
--- $\sum_{j}\frac{1}{4}\Delta_{1}(1-\gamma)(\sigma_{j-1}^{x}\sigma_{j}^{x}\sigma_{j+1}^{z}+\sigma_{j-1}^{z}\sigma_{j}^{x}\sigma_{j+1}^{x})$ (S12)
-Using $\langle\sigma_{j}^{x}\rangle_{0}=\sin{\theta}$ ) and $\langle\sigma_{j}^{z}\rangle_{0}=\cos{\theta}$ ). We find:
-$\langle[H,[H,Q]]\rangle_{0}=\frac{1}{2}(1-\gamma)((1-\gamma)L\cos{\theta}+L\cos{\theta}\sin^{2}{\theta})-\frac{1}{2}\Delta_{1}(1-\gamma)L\cos{\theta}\sin^{2}{\theta}$ (S13)
-Next, we evaluate the term $[H,[H,Q^{2}]]$. Expanding this expression, we obtain:
-$[H,[H,Q^{2}]]$ $=$ $[H,Q[H,Q]]+[H,[H,Q]Q]$ (S14)
-$=$ $Q[H,[H,Q]]+[H,[H,Q]]Q+2[H,Q]^{2}$ (S15)
-The expectation value of the third term is:
-$\langle[H,Q]^{2}\rangle_{0}$ $=$ $-\frac{1}{4}(1-\gamma)^{2}\langle\sum_{i}(\sigma_{i}^{x}\sigma_{i+1}^{y}+\sigma_{i}^{y}\sigma_{i+1}^{x})\sum_{j}(\sigma_{j}^{x}\sigma_{j+1}^{y}+\sigma_{j}^{y}\sigma_{j+1}^{x})\rangle_{0}$ (S16)
-$=$ $-\frac{1}{2}(1-\gamma)^{2}L(1+\sin^{2}{\theta})$ (S17)
-The first term, $\langle Q[H,[H,Q]]\rangle_{0}$, is computed as:
-$\langle Q[H,[H,Q]]\rangle_{0}$ $=$ $\langle\frac{1}{2}(1-\gamma)^{2}\sum_{i,j}\sigma_{i}^{z}\sigma_{j}^{z}+\frac{1}{2}(1-\gamma)\sum_{i,j}\sigma_{i}^{z}\sigma_{j}^{z}(\sigma_{j-1}^{x}\sigma_{j+1}^{x}-\gamma\sigma_{j-1}^{y}\sigma_{j+1}^{y})\rangle_{0}$ (S18)
--- $\langle\sum_{i,j}\frac{1}{4}\Delta_{1}(1-\gamma)(\sigma_{i}^{z}\sigma_{j-1}^{x}\sigma_{j}^{x}\sigma_{j+1}^{z}+\sigma_{i}^{z}\sigma_{j-1}^{z}\sigma_{j}^{x}\sigma_{j+1}^{x})\rangle_{0}$ (S19)
-$=$ $\frac{1}{2}(1-\gamma)^{2}(L+L(L-1)\cos^{2}{\theta})+\frac{1}{2}(1-\gamma)L(L-3)\sin^{2}{\theta}\cos^{2}{\theta}+\frac{1}{2}(1-\gamma)L\sin^{2}{\theta}$ (S20)
--- $\frac{1}{2}\Delta_{1}(1-\gamma)(L(L-3)\sin^{2}{\theta}\cos^{2}{\theta}+L\sin^{2}{\theta})$ (S21)
-Here, we have used $\langle Q^{2}\rangle_{0}=L+L(L-1)\cos^{2}{\theta}$.
-The second term in Eq.
-S18 has two contributions from $i=j$ and $i\neq j$, $j-1$, and $j+1$, while the third term vanishes due to $\langle\sigma_{j}^{y}\rangle_{0}=0$.
-Moreover, $\langle Q[H,[H,Q]]\rangle_{0}=\langle[H,[H,Q]]Q\rangle_{0}$.
-Collecting all terms in Eq.
-S15, the first term in Eq.
-S9 becomes
-$-\langle[H,[H,Q^{2}]]\rangle_{0}$ $=$ $-(1-\gamma)^{2}(L+L(L-1)\cos^{2}{\theta})-(1-\gamma)L(L-3)\sin^{2}{\theta}\cos^{2}{\theta}-(1-\gamma)L\sin^{2}{\theta}$ (S22)
-$+$ $\Delta_{1}(1-\gamma)(L(L-3)\sin^{2}{\theta}\cos^{2}{\theta}+L\sin^{2}{\theta})+(1-\gamma)^{2}L(1+\sin^{2}{\theta})$ (S23)
-$=$ $-(1-\gamma)^{2}(L^{2}\cos^{2}{\theta}-L)-(1-\gamma)(L^{2}-3L)\sin^{2}{\theta}\cos^{2}{\theta}-(1-\gamma)L\sin^{2}{\theta}$ (S24)
-$+$ $\Delta_{1}(1-\gamma)(L^{2}-3L)\sin^{2}{\theta}\cos^{2}{\theta}+\Delta_{1}(1-\gamma)L\sin^{2}{\theta}$ (S25)
-The second term in Eq. S9 is computed as:
-[…]
-$=$ $L^{2}(1-\gamma)^{2}\cos^{2}{\theta}+L^{2}(1-\gamma)\cos^{2}{\theta}\sin^{2}{\theta}-\Delta_{1}(1-\gamma)L^{2}\cos^{2}{\theta}\sin^{2}{\theta}$ (S27)
-Combining all terms, the second derivative of the charge variance is:
-$\left.\frac{d^{2}\sigma_{Q}^{2}}{dt^{2}}\right|_{0}$ $=$ $-\langle[H,[H,Q^{2}]]\rangle_{0}+2\langle Q\rangle_{0}\langle[H,[H,Q]]\rangle_{0}$ (S28)
-$=$ $-(1-\gamma)^{2}(L^{2}\cos^{2}{\theta}-L))-(1-\gamma)(L^{2}-3L)\sin^{2}{\theta}\cos^{2}{\theta}-(1-\gamma)L\sin^{2}{\theta}$ (S29)
-$+$ $\Delta_{1}(1-\gamma)(L^{2}-3L)\sin^{2}{\theta}\cos^{2}{\theta}+\Delta_{1}(1-\gamma)L\sin^{2}{\theta}$ (S30)
-$+$ $L^{2}(1-\gamma)^{2}\cos^{2}{\theta}+L^{2}(1-\gamma)\cos^{2}{\theta}\sin^{2}{\theta}-\Delta_{1}(1-\gamma)L^{2}\cos^{2}{\theta}\sin^{2}{\theta}$ (S31)
-$=$ $L(1-\gamma)^{2}+3L(1-\gamma)\sin^{2}{\theta}\cos^{2}{\theta}-L(1-\gamma)\sin^{2}{\theta}$ (S32)
--- $3\Delta_{1}L(1-\gamma)\sin^{2}{\theta}\cos^{2}{\theta}+\Delta_{1}L(1-\gamma)\sin^{2}{\theta}$ (S33)
-The early-time growth of the charge variance for different values of $\theta$ is then approximated by:
-$\sigma_{Q}^{2}(t)/L=\sin^{2}{\theta}+\frac{t^{2}}{2}(1-\gamma)\left(1-\gamma+(1-\Delta_{1})(3\sin^{2}\theta\cos^{2}\theta-\sin^{2}\theta\right))$ (S34)
-Here, the initial charge variance, $\sigma_{Q}^{2}(0)$, is replaced with $L(1-\cos^{2}{\theta})$.
-We can now compare this analytical result with numerical simulations at early times.
-We find that the charge variance obtained from both methods is in good agreement at early times, as confirmed in Fig.
-S23.
-Figure S23: Comparison of the early-time behavior of the charge variance, $\sigma_{Q}^{2}$, between analytical solutions (dashed lines) and numerical simulations (solid lines).
-Results are obtained from tilted ferromagnetic states under the Hamiltonian $H_{1}$ with $L=12$.
-Panel (a): $\gamma=0.7$.
-Panel (b): $\gamma=0.6$.